### v1

Abstract
We study polarization in a probabilistic voting model with aggregate shocks and a decreasing marginal utility from office rents. In equilibrium, parties offer different policies, despite being rent-motivated and ex-ante identical from the point of view of voters. When candidates compete on a single policy issue, parties’ equilibrium payoffs increase in voter polarization, even when the change is driven by the supporters of the opposite party becoming more extreme. With multiple policy issues, parties benefit if the society is split into two factions and the ideological cohesion within such factions increases. We connect our results to empirical evidence on polarizing political communication, party identity, and zero-sum thinking, and find that polarization could be reduced by intervening on the electoral rule.
JEL Codes: D72, D74, D81, D90.
Keywords: polarization; conflict; elections; zero-sum thinking; identity.
1 Introduction
Americans are divided along ideological and partisan lines more than ever in the past. This polarization can be observed both in the electorate and between political parties. On the one hand, US voters are more and more clustered in two factions taking opposite stances on a wide array of policy issues. [^1] https://www.pewresearch.org/politics/interactives/political-polarization-1994-2017/ provides a user-friendly interactive graph that illustrates this point, also documented by Gentzkow, (2016), Bonomi et al., (2021) and Bertrand and Kamenica, (2023). Political divisions, polarization, and gridlock are perceived by many Americans as some of the most important problems affecting their country (Pew Research Center,, 2023). On the other hand, politicians from opposite parties appear more and more different and polarized in their speeches (Gentzkow et al.,, 2019) and vote choices McCarty, (2016).
As a result of these recent developments, the issue of polarization and its determinants is attracting more and more attention from social scientists. [^2] According to Google Ngrams, the share of books containing the terms “voter polarization” more than tripled between $1998$ and $2018$. The share of books containing the phrase “party polarization” was more than seven times as large in 2018 as in 1998. A first set of explanations stresses the demand-side determinants of political conflict, including the occurrence of structural economic shocks (Autor et al.,, 2020) and their interaction with social identity in exacerbating existing divisions in the society (Bonomi et al.,, 2021), as well as the role of technological innovations facilitating the creation of online echo-chambers resulting in belief distortions and polarization (Bowen et al.,, 2023). A second set of explanations stresses the role of political supply-side actors and information sources in reinforcing voters’ conflict. Indeed, exposure to slanted news has been found to influence voters’ political views (Martin and Yurukoglu,, 2017; DellaVigna and Kaplan,, 2007), and media outlets have become increasingly polarized over time (Kim et al.,, 2022). Recent empirical evidence also suggests that parties’ propaganda interacts with the partisanship of the audience in a way that can increase issue polarization between opposite party supporters (Flores et al.,, 2022; Cohen,, 2003), often because of a strong backlash effect among misaligned voters (Feddersen and Adams,, 2022; Ciuk,, 2018; Goren et al.,, 2009). [^3] Interestingly, party choices can polarize the electorate in less active ways. For instance, the entrance of extremist politicians in parliament can generate voter polarization through a mixture of legitimization and backlash effects (Bischof and Wagner,, 2019).
The relation between voter polarization and party polarization (hereafter, party divergence) and the reason why politicians seem to communicate in ways that polarize voters, potentially triggering a backlash among disagreeing social groups, remain open questions. The goal of this paper is to contribute to the literature addressing these questions.
Our first contribution is to advance a novel explanation for party divergence in two-party systems. The issue of why competing candidates would offer different policies in an electoral competition equilibrium is a non-trivial one, especially if they are motivated exclusively by political rents, as in the seminal work by Downs, (1957) and common in the economic literature (Lindbeck and Weibull,, 1987; Persson and Tabellini,, 2002). Typically, policy divergence requires parties to be somewhat asymmetric to start with – either because they are driven by different policy goals (Wittman,, 1977; Hansson and Stuart,, 1984; Lindbeck and Weibull,, 1993), followed by different social groups (e.g., Glaeser et al.,, 2005; Gennaioli and Tabellini,, 2023), or because very different voter groups participate in the respective primaries (e.g., Brady et al.,, 2007; Hirano et al.,, 2009). These are realistic assumptions, although they leave the question of the origin of these asymmetries open.
We address this aspect, by presenting a probabilistic voting model where party divergence occurs in equilibrium even if parties are ex-ante identical in all model primitives. As standard, party payoffs depend positively on the share of office rents they can secure after the election, which are increasing in their vote share (e.g., Lindbeck and Weibull,, 1987; Persson and Tabellini,, 2002; Glaeser et al.,, 2005). We depart from the standard model by assuming that politicians’ marginal utility from rents is decreasing, for instance, because rents are their major source of income and therefore consumption. We find that the presence of aggregate popularity shocks pushes different politicians to cater to different voter groups in equilibrium. Why do parties diverge? Divergence allows parties to insure against the risk of becoming politically irrelevant, as targeting the demands of different parts of the population allows them to build a core voter base and secure at least some office rents.
We then build on this first result and unveil a novel link between voter polarization, party divergence, and parties’ equilibrium payoffs. Focusing on competition over a single policy issue, say government spending, we show that both parties’ equilibrium payoffs increase when the policy views of those to the left of the median voter become further left-wing while right-wing voters move further to the right of the political spectrum. When the political divisions among voters sharpen, the equilibrium core supporters of the left-wing and right-wing parties have very different policy demands, and parties can increase their degree of differentiation and better appeal to their core voters. The same rationale yields a somewhat unexpected, but relevant, implication: each candidate always benefits from alienating at least some of the competitor’s supporters, with communication strategies that make them more extreme, even if this reduces the support from these voters in equilibrium. By alienating left-wing voters, the right-wing politician effectively pushes her competitor to propose an even more ambitious redistribution program, making her own fiscally conservative stances relatively more appealing to the right wing of the electorate. These results suggest a new perspective on the role of political parties in fueling voter conflict: under some circumstances, triggering a “targeted backlash” might advantage politicians.
Next, we extend the comparative statics about voter conflict to a multidimensional policy space. The core intuition of the single-policy case carries over if politicians run on multiple policies: increasing voter conflict can benefit both parties. However, in a multidimensional policy arena, a divided electorate is not necessarily one that is highly polarized over specific issues. Instead, we show that politicians are better off, in equilibrium, when the society becomes further fractured into two blocks of citizens that are very distant from each other in all their policy demands, much as described at the beginning of this section. Indeed, our results suggest that parties are better off when voters’ policy views are correlated rather than independent across different policy domains, as when opposite voter groups are highly ideologically cohesive.
Our main theoretical results highlight a novel incentive to exacerbate societal divisions. In the conclusive sections of the paper we discuss examples of how this might occur, relating the main themes of the previous sections to the growing literature on zero-sum thinking (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), common interest vs conflict of interest policy issues (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023), and identity politics (e.g. Shayo,, 2009; Bonomi et al.,, 2021; Gennaioli and Tabellini,, 2023). First, our discussion suggests that politicians might benefit if voters overestimate the zero-sum component of politics and perceive conflict of interest political issues as more salient than common interest policies, as this causes disagreement between opposite-minded voters to intensify.
Similarly, we predict the political supply to over-provide information regarding conflict of interest policy issues – those where different policies benefit different voters – and under-provide information on common interest policies, causing a loss of (utilitarian) welfare. On the one hand, as voters learn which policy accommodates their particular interests, they become more divided. This allows parties to diverge more in equilibrium but also creates more uncertainty about which platform will be implemented, hurting (risk-averse) voters. On the other hand, information about common interest policies is valuable to voters because it forces politicians to converge on the socially optimal proposal, but has no value to the policy-motivated candidates. [^4] Indeed, if voters had heterogeneous priors or idiosyncratic private information, public information about common interest issues would bring them closer together, decreasing the equilibrium expected payoff of politicians.
Our last application concerns party identity, modeled building on (Bonomi et al.,, 2021). We show that prompting voters to identify with the party they are more likely to vote for is not only valuable but also “efficient” for politicians, in that it distorts all voters’ bliss points in the direction of the fastest payoff improvement for candidates. These results complement those of Gennaioli and Tabellini, (2023), who show how competing parties might leverage other types of social identities for their electoral advantage. We also illustrate that politicians’ differentiation incentives can interact with voters’ party identity, with a feedback loop between party divergence and voter conflict that can generate a growth in polarization over time. Finally, we discuss how changing some features of the electoral rule might curb party divergence and the incentives to divide voters.
The main results of this paper are related, either from a technical or high-level perspective to two sets of recent contributions to the political economy literature. From a technical perspective, Zakharov and Sorokin, (2014) show that if voters have heterogeneous preferences both over parties and policies and popularity shocks are idiosyncratic, equilibrium platform convergence requires the spatial competition game to be a constant sum, while Sorokin and Zakharov, (2018) demonstrate that, even in such a setup, party convergence always occurs if the number of voters is large. In our model, with aggregate shocks, ex-ante identical parties, and constant aggregate office rents, both results do not hold anymore: on the one hand, platform convergence is expected in many cases in which the game is not a constant sum, as discussed in section 2; on the other hand, with a decreasing marginal utility from rents, policy divergence occurs regardless of the number of voters, as long as voters are not perfectly homogeneous in their policy preferences. [^5] Sorokin and Zakharov, (2018) point out that in the presence of aggregate shocks party payoffs might still matter, although their discussion applies to the case where voters have a preference for parties on top of policies, which facilitates divergence results. We believe that focusing on the case where parties are ex-ante identical, both in their motivation and from the perspective of voters, is per se interesting, as it allows focusing on polarization forces that do not rely on path dependence.
From a high-level perspective, our paper is related to Glaeser et al., (2005) and Gennaioli and Tabellini, (2023), who assume that parties are historically connected to different social groups, and show how they have the incentive to render their respective voter groups more extreme. Their argument hinges on candidates’ ability to target voters or, equivalently, leverage voters’ attention asymmetries: catalyzing their own supporters’ extremism pays off provided that it does not equally alienate those on the opposite side of the political spectrum. In this paper, we provide a complementary view, where what matters the most to parties is exacerbating the divisions present in the electorate, regardless of whether this is achieved by fomenting those voters who will become their core supporters or alienating the ones that will turn to their competitors. [^6] In our comparative statics, we compare equilibrium payoffs under different voter distributions, so that both parties are allowed to react to changes in the voter distribution by adjusting their platforms. With fixed platforms, it is possible to show that each party could benefit from propaganda that increases voter disagreement, even when its main (but not only) effect was alienating the core supporters of the competitor.
The paper is organized as follows. In section 2 we describe the general model features and present our equilibrium divergence result. In sections 3 and 4 we focus on a simple specification of the model and present our comparative statics on polarization and voter conflict for the unidimensional and multidimensional cases respectively. In section 5 and 6 we discuss applications and the relation between polarization and the institutional framework. Finally, in section 7 we conclude.
2 Model
Two parties, $A$ and $B$ compete in an election. The electorate is described by a distribution of voter types $(x_{i},s_{i})_{i\in I}$, where $i=1,...,N$ indexes a type, $x_{i}\in\mathbb{R}^{K}$ is the vector of policy bliss points of voter type $i$, and $s_{i}$ is the share of voters of type $i$. The electorate is diverse, that is, the support of the voter distribution contains at least two types. The utility of voter $i=1,...,N$ from policy $x\in\mathbb{R}^{K}$ is $v(x,x_{i})=g(x^{i})+\sum_{k\leq K}h^{k}(|x^{k}-x^{k}_{i}|)$, with $g:\mathbb{R}^{K}\to\mathbb{R}$ and $h^{k}:\mathbb{R}\to\mathbb{R}$ for each $k=1,...,K$. Each $h^{k}$ is strictly concave and decreasing in $|x^{k}-x^{k}_{i}|$.
The candidates of each party simultaneously announce their platforms $x_{A},x_{B}\in\mathbb{R}^{K}$. After the platforms are announced, but before citizens cast their vote, a popularity shock in favor of the candidate of party $B$, $\epsilon\in\mathbb{R}$, is realized. The shock has a continuous cumulative distribution $F$, and captures random events that influence voters’ relative assessment of party $B$, including, for instance, new information about the state of the economy, political scandals, or the effect of a candidate’s personality traits and characteristics on their appeal to voters. We assume that voters vote sincerely, taking into account $x_{A}$, $x_{B}$, and the popularity shock $\epsilon$. In particular, for each $x_{A},x_{B}$, let $\Delta_{i}(x_{A},x_{B})=v(x_{A},x_{i})-v(x_{B},x_{i})$. Each voter of type $i$ will vote for party $A$ if and only if $\Delta_{i}(x_{A},x_{B})\geq\epsilon$, which holds with probability $\pi_{i}(x_{A},x_{B})=F(\Delta_{i}(x_{A},x_{B}))$. As standard in probabilistic voting models, we let the support of $F$ be wide, so that cases where $\pi_{i}\in\{0,1\}$ can be ignored in the equilibrium analysis. [^7] Formally, for $x_{\lor}=(\max_{i\in I}{x^{k}_{i}})_{k\leq K}$ and $x_{\land}=(\min_{i\in I}{x^{k}_{i}})_{k\leq K}$, we assume that both $v(x_{\land},x_{\lor})-v(x_{\lor},x_{\lor})$ and $v(x_{\lor},x_{\lor})-v(x_{\land},x_{\lor})$ are in the support of $F$.
Politicians are exclusively motivated by office rents, which we assume to depend positively on the share of total votes that are cast in their favor. In particular, let $R:[0,1]\to\mathbb{R}$ be the rent allocation function, mapping the vote share $s$ of a candidate to rents $R(s)$ for her party. We make the following assumptions about $R$.
Assumption 1
The rent allocation function $R:[0,1]\to\mathbb{R}$ has the following two properties: (i) $R$ is strictly increasing on $[0,1]$; (ii) there is some constant $\bar{R}\in\mathbb{R}$, such that $R(s)+R(1-s)=\bar{R}$ for each $s\in[0,1]$.
Part (i) requires that rents increase in the number of votes. [^8] Proposition 2.1 would generalize to the case where $R$ is weakly increasing on $[0,\frac{1}{2}]$ and not constant on $[0,\frac{1}{2})$. In that case, party divergence would obtain for some distributions of voter types. We do not assume continuity: this allows, for instance, for jumps in the office rents as a party’s vote share meets institutional thresholds that grant agenda-setting power, the power to block legislation or entitlement to electoral reimbursements. In particular, the general setup allows for a party to gain a sizable share of power (and political rents) once it obtains the absolute majority of votes. An alternative interpretation of this setup (proposed, e.g., by Glaeser et al.,, 2005), is that voters are making a campaign contribution decision rather than a voting decision, so that parties’ rents naturally grow with their supporters. [^9] Under this interpretation, the size $s_{i}$ of voter group $i$ would already account for the heterogeneity in voters’ budget devoted to political contributions. Part (ii) imposes the restriction that aggregate rents are constant. Our second main assumption is that parties’ utility from rents is increasing and concave. Formally, let the payoff function of each party be $U:\mathbb{R}\to\mathbb{R}$. The following holds.
Assumption 2
The utility function $U:\mathbb{R}\to\mathbb{R}$ is such that $U^{\prime}>0$ and $U^{\prime\prime}<0$.
We think of the office rents as having direct or indirect monetary value to politicians, so that the concavity of $U$ follows from a decreasing marginal utility from consumption. An alternative interpretation is that risk-neutral politicians might use the rents to maintain their support through transfers that increase their constituents’ (concave) consumption utility, or by investments in a propaganda technology with convex costs.
Let $\nu:[0,1]\to\mathbb{R}$, $\nu=U\circ R$, be the function mapping a party’s vote share to its rents, and without any loss of generality, let payoffs be normalized so that $\nu(1)-\nu(0)=1$. The following holds when assumptions 1 and 2 are combined.
Remark 1
For each $s,s^{\prime}\in[0,\frac{1}{2}]$, with $s^{\prime}>s$, it holds that $\nu(s^{\prime})-\nu(s)>\nu(1-s)-\nu(1-s^{\prime})$.
In other words, our assumptions imply that winning an additional group of voters is worth more to the minority party than to the majority party.
Having outlined the characteristics of the model, we can now turn to the equilibrium analysis of the party competition game.
2.1 Party Divergence
[…]
Our first result is that in equilibrium the two candidates run with different platforms with positive probability.
Proposition 2.1 (Party Divergence)
In every pure strategy equilibrium, candidates offer different policy platforms.
The intuition for the result is as follows. If parties mimic each other’s platforms, candidates look ex-ante identical from the voters’ perspective, so the election will be entirely decided by random factors (i.e., the realization of the popularity shock $\epsilon$). In contrast, by offering a platform different from the one of their opponent, a candidate can better meet the policy requests of some voter types, “securing” at least some votes and power rents. This differentiation might lead to a lower expected vote share, but it also reduces the uncertainty in the electoral outcome, acting as insurance, and increasing the candidate’s expected utility.
An analogous intuition would suggest that party convergence would be reinstated in equilibrium if the utility from rents was convex, $U^{\prime\prime}>0$. It is easy to show that this is indeed the case. [^11] More precisely, one can show that if candidates converge when each candidate $p$ maximizes her expected rents $R(s_{p})$, then if $U^{\prime\prime}>0$, they also converge when maximizing $U(R(s_{p}))$. To see this, let strategy profile $x^{\star}=(x,x)$ be any (pure strategy) equilibrium when parties are expected-rent maximizers. Since parties offer the same policy, $p$ obtains $\bar{R}$ with probability $\frac{1}{2}$ and 00 otherwise, with an expected rent $\frac{\bar{R}}{2}$. Assume that now candidate $p$ maximizes $U(R(s_{p}))$, with $U^{\prime\prime}>0$ and consider $x^{\star}$ as candidate equilibrium. If $p$ deviates, she has an expected rent of at most $\frac{\bar{R}}{2}$, with their rents following some distribution $G$, with support different from $\{0,\bar{R}\}$. Now let $H$ be mean preserving spread of $G$, supported only on $\{0,\bar{R}\}$. Because $U^{\prime\prime}>0$, candidate $p$ is better off under $G$ than $H$. But $G$ is either equal to or first-order dominated by the rent distribution induced by $x^{\star}$, so the deviation cannot be profitable. It turns out that the party divergence result of proposition 2.1 holds for all voter type distributions and voter utilities satisfying our assumptions if and only if $\nu(s)+\nu(1-s)>\nu(1)+\nu(0)$ for each $s\in(0,1)$, a condition weaker than remark 1, but not satisfied when $U^{\prime\prime}>0$.
In the next sections, we consider a particularly tractable case of the above general formulation. These tractability assumptions on voters’ utility and shocks are common in the literature (see, e.g., Gennaioli and Tabellini,, 2023; Glaeser et al.,, 2005). In our setup, they come with the benefit of characterizing the political equilibrium and shedding light on one of this paper’s most interesting and novel results: parties’ incentives to polarize voters.
3 Unidimensional Conflict
We start by analyzing the case where $K=1$, so that parties compete on a single policy dimension. [^12] In this case, dimension indexes $k$ are redundant, and omitted. Let $\epsilon\sim U(-\phi,\phi$), so that the shock is uniformly distributed, and assume that $h^{k}(|x-x_{i}|)=-\alpha(x^{k}-x^{k}_{i})^{2}$, where $\alpha>0$ captures how important policies are for individual payoffs and, without loss of generality, we set $\alpha=1$. With one policy dimension, we can think of voter heterogeneous preferences as capturing the classical economy-based conflict of interest between right-wing and left-wing voters, as in the following simple example.
Redistributive Conflict
Income is distributed in the society following $(y_{i},s_{i})_{i\in I}$ where $y_{i}\in\mathbb{R}$ is the income of income type $i$ and $s_{i}$ is the proportion of such a type in the population, and the mean income is $\bar{y}=1$. The utility of a voter with income type $y_{i}$ is given by $u(y_{i},\tau)=(1-\tau)y_{i}+G(\tau)$, where $(1-\tau)y_{i}$ is the utility from income, net of an income tax rate $\tau$, while $G(\tau)=\psi\tau-\frac{c}{2}\tau^{2}$ is the utility from public good provision, net of inefficiencies originated from taxation, where $\psi>0$ and $c>0$, both large. [^13] If $\psi$ and $c$ are large enough, then the desired tax rate of each individual is non-negative and lies in $[0,1]$. It is easy to see that the taxation bliss point (i.e. ideal policy) of income type $y_{i}$ is $\tau_{i}=\frac{\psi-y_{i}}{c}$, so that higher income types demand lower taxes. Note that the utility function of voter $i$ can be rewritten as $y_{i}+\frac{(\psi-y_{i})^{2}}{2c}-\frac{c}{2}(\tau-\tau_{i})^{2}$. Hence, this example satisfies the assumptions of our model for $x=\tau$, $x_{i}=\tau_{i}$, and $g(x_{i})=\psi-cx_{i}+\frac{c}{2}x_{i}^{2}$.
First, we show that the game admits a pure strategy equilibrium, and there are exactly two platforms that can be offered in such an equilibrium. One policy lies to the left of the risk-neutral equilibrium policy $x^{RN}$. The other lies to the right of such policy. For the next propositions, assume that voters are ordered so that $x_{1}\leq x_{2}\leq...\leq x_{N}$, and let $x_{m}$ be the median voter type. If the median voter does not exist, define $x_{m}$ to be the average of the two “most median” voter types.
Proposition 3.1 (Unidimensional Equilibrium)
The game admits a pure strategy equilibrium. Moreover, there exist two platforms, $\underaccent{\bar}{x},\bar{x}\in\mathbb{R}$ with $x_{1}<\underaccent{\bar}{x}<\bar{x}<x_{N}$, such that, in every pure strategy equilibrium, one candidate offers $\underaccent{\bar}{x}$ while the other candidate offers $\bar{x}$. It has
$\underaccent{\bar}{x}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i},\quad\quad\quad\quad\bar{x}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$
where $w_{i},\omega_{i}\in(0,1)$, $\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=\sum_{i=1}^{N}\bar{w}_{i}=1$, and $\bar{w}_{i}>(<)\text{ }\underaccent{\bar}{w}^{i}$ if $x_{i}>(<)\text{ }x_{m}$, where $x_{m}$ is the median voter type. For each $i$, the equilibrium weights $\bar{w}_{i}$ and $\underaccent{\bar}{w}_{i}$ that parties attribute to voter group $i$ are proportional to their payoff increase from winning the votes of such group. Finally, the equilibrium expected payoff of each party is proportional to the (squared) distance between the two equilibrium platforms $(\bar{x}-\underaccent{\bar}{x})^{2}$, up to a constant term.
Remark 2
If the voter distribution is symmetric or the share of political rents is proportional to the vote share, then $\underaccent{\bar}{x}<x^{o}<\bar{x}$, where $x^{o}$ maximizes aggregate voter welfare.
Using our simple model of redistributive conflict to interpret the results, proposition 2 tells us that in equilibrium one party is more left-wing than the other, promising a higher tax rate if elected. Each party will have its “core voters:” The very poor (rich) are expected to be relatively more likely to support the party that offers more (less) redistribution. Interestingly, the core voters are endogenous: the result emerges from the decreasing marginal utility from rents, rather than from asymmetries in the primitive of the model. Naturally, party divergence is welfare-detrimental if, as in the cases of remark 2, the risk-neutral equilibrium policy maximizes voters’ (utilitarian) welfare: when parties offer different platforms, a voter’s benefits from the potential election of their favorite candidate are more than compensated by the risk of the electoral success of a candidate targeting those on the other end of the political spectrum. Finally, parties are better off the more they can appear different in equilibrium: the further apart their equilibrium platforms, the less likely their core voters will be to vote for the other party. This last result sets the grounds for our findings on voter polarization, presented in the next section.
3.1 Voter Alienation and Polarization
Without loss of generality, in this section, we assume that $x_{A}=\bar{x}$ and $x_{B}=\underaccent{\bar}{x}$. Before stating the next results formally, it is useful to introduce some notation and definitions that will ease the analysis. First, let $\mathbf{(x,s)}=(x_{i},s_{i})_{i\in I}$ represent the voter type distribution. We can write the equilibrium platforms and payoffs as a function of such distribution. In particular we denote by $\bar{x}\mathbf{(x,s)}$ and $\underaccent{\bar}{x}\mathbf{(x,s)}$ the equilibrium platforms as functions of $\mathbf{(x,s)}$ and let $\hat{V}_{p}(\mathbf{x},\mathbf{s})$ be the corresponding equilibrium expected payoff of party $p$. Similarly, for $p=A,B$ and $i\in I$ denote by $\pi^{p}_{i}(\mathbf{x,s})$ the probability that group $i$ will vote for $p$ when the equilibrium policies are consistent with $\mathbf{(x,s)}$.
Definition 1
Fix a voter distribution $\mathbf{(x,s)}$. We say that candidate $p\in\{A,B\}$ benefits from attracting voter group $i\in I$ if $\frac{\partial\pi^{p}_{i}}{\partial x_{i}}\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}>0$, while we say that she benefits from alienating voter group $i\in I$ if $\frac{\partial\pi^{p}_{i}}{\partial x_{i}}\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}<0$.
In other words, a right-wing candidate benefits from attracting voter type $i$ if a change in $i$’s views on redistribution pushes her to support the political right and also increases the equilibrium payoff of the right-wing party. Conversely, the right-wing party benefits from alienating voter type $i$ if an increase in $i$’s support for redistribution makes the right-wing party better off in equilibrium, despite increasing left-wing support from group $i$.
At this point, one could be skeptical that a candidate ever benefits from alienating a group of voters, even if it means that the rest of the voter distribution remains unchanged and that group of voters will be less likely to support the party in the new equilibrium. Surprisingly, our next result suggests that each party always has the incentive to alienate at least one voter group: the most extremist supporters of the rival candidate.
Proposition 3.2
Let $(\mathbf{x},\mathbf{s})$ be any voter distribution and $x_{m}$ be the median voter type. Party $A$ benefits from attracting group $i$ if $x_{i}>\max\{\underaccent{\bar}{x}(\mathbf{x},\mathbf{s}),x_{m}\}$ and alienating group $i$ if $x_{i}<\min\{\underaccent{\bar}{x}(\mathbf{x},\mathbf{s}),x_{m}\}$. Party $B$ benefits from alienating group $i$ if $x_{i}>\max\{\bar{x}(\mathbf{x},\mathbf{s}),x_{m}\}$ and attracting group $i$ if $x_{i}<\min\{\bar{x}(\mathbf{x},\mathbf{s}),x_{m}\}$.
Remark 3
Let $x_{m}\notin\{x_{1},x_{N}\}$. Party $A$ always benefits from attracting group $N$ and alienating group $1$. Party $B$ always benefits from attracting group $1$ and alienating group $N$.
A closer look at how the equilibrium utility of party $B$ changes when $x_{N}$ becomes marginally more extreme can help understand the result. Let $V_{p}(x_{A},x_{B})$ be the expected utility of party $B$ when platforms $x_{A}$ and $x_{B}$ are offered in equilibrium. Since we are assuming that $x_{A}=\bar{x}$ and $x_{B}=\underaccent{\bar}{x}$, voters are sorted in ascending order in terms of relative preference for party $A$. It has, [^14] While the objective function of each party is not everywhere differentiable, we show in the appendix that the equilibrium platforms satisfy the FOC of the maximization problem.
$\frac{\partial\hat{V}_{B}(\mathbf{x},\mathbf{s})}{\partial x_{N}}=\underbrace{\frac{\partial V_{B}(\bar{x}(\mathbf{x},\mathbf{s}),\underaccent{\bar}{x}(\mathbf{x},\mathbf{s}))}{\partial x_{N}}}_{\begin{subarray}{c}\text{direct alienation effect}\\ <0\end{subarray}}+\underbrace{\frac{\partial V_{B}(\bar{x}(\mathbf{x},\mathbf{s}),\underaccent{\bar}{x}(\mathbf{x},\mathbf{s}))}{\partial\bar{x}}\frac{\partial\bar{x}(\mathbf{x},\mathbf{s})}{\partial x_{N}}}_{\begin{subarray}{c}\text{effect of party A}\\ \text{becoming more extreme}\\ >0\end{subarray}}.$ (1)
The first term on the right-end-side captures the direct effect of group $N$ becoming relatively further apart from party $B$, and more unlikely to vote for its candidate. This effect is naturally negative, but its payoff relevance is attenuated since group $N$ can only vote for $B$ if $B$ already has the support of all other groups, that is, if $B$ can extract high rents even without the support of $N$. The second term is the indirect effect that an increase in $x_{N}$ has on the payoff of $B$ through the readjustment it induces in $\bar{x}$, the equilibrium platform of party $A$. Note that $\frac{\partial\bar{x}(\mathbf{x},\mathbf{s})}{\partial x_{N}}=\bar{w}_{N}>0$, which means that party $A$ moves its platform in the same direction of $x_{N}$, and away from the core supporters of $B$ [^15] Note that also party $B$’s equilibrium platform accommodates the change in $x_{N}$, that is, $\frac{\partial\underaccent{\bar}{x}(\mathbf{x},\mathbf{s})}{\partial x_{N}}=\omega_{N}>0$. However, because $x_{N}>x_{m}$, we have $w_{N}>\omega_{N}$ by proposition 3.1, so party $B$ is less responsive to the change in $x_{N}$ than party $A$. This effect drops out of equation 1 by the envelope theorem. As a result of this change, party $B$ will be more likely to secure the vote of its core supporters and the associated political rents. It can be shown that this second effect dominates the others.
To sum up the intuition: by alienating the more extreme left-wing voters, a right-wing candidate can successfully push the left to propose equilibrium platforms that are further apart from the demands of far-right voters, further securing their support. The same reasoning applies to the right-wing candidate.
In general, it can be shown that for each voter distribution $(\mathbf{x},\mathbf{s})$, voter group $i$ and party $p$ it holds,
$x_{i}>(<)\text{ }x_{m}\implies\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}>(<)\text{ }0,$
which highlights a common interest for the competing candidates. Both parties are better off, in equilibrium, when voters whose bliss point is to the left of the median move further to the left, and those to the right move further to the right. [^16] Note that this does not mean that the party offering $\underaccent{\bar}{x}$ benefits from alienating all those groups that are to the right of the median voter. If the voter distribution is sufficiently skewed, $x_{B}=\underaccent{\bar}{x}$ could be greater than $x_{m}$, and some voter groups to the right of the median voter might become more likely to vote party $B$ when their ideal policies move further to the right. This result holds beyond marginal changes: both political parties benefit when the distribution of voters is more spread out, and in particular, the disagreement between the top $50\%$ and the bottom $50\%$ of voters increases. Before formally stating the result, it is useful to introduce the notion of spread. For a given distribution of voters $\mathbf{(x,s)}$, denote by $(\mathbf{x}_{L},\mathbf{s}_{L})$ and $(\mathbf{x}_{R},\mathbf{s}_{R})$ the distribution conditional on voters belonging to, respectively, the bottom and the top $50\%$ of $\mathbf{(x,s)}$.
Definition 2
Distribution $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ is a spread of distribution $\mathbf{(x,s)}$ if $(\mathbf{x}_{L},\mathbf{s}_{L})$ first-order stochastically dominates $(\mathbf{x}^{\prime}_{L},\mathbf{s}^{\prime}_{L})$ and $(\mathbf{x}^{\prime}_{R},\mathbf{s}^{\prime}_{R})$ first-order stochastically dominates $(\mathbf{x}_{R},\mathbf{s}_{R})$, with at least one of the two relations holding strictly.
If a voter distribution is a spread of another, then voters to the left (right) of the median voter are more likely to hold far-left (far-right) ideal policies in the former than in the latter distribution, so that conflict between opposite extremists is exacerbated. We find that politicians are better off when the voter distribution spreads.
Proposition 3.3
Parties are better off, in equilibrium, when the conflict between those to the left and those to the right of the median voter increases. More formally, let $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ and $\mathbf{(x,s)}$ be two voter distributions, and let $(\mathbf{x}^{\prime}-\mathbf{x}^{\prime}_{m},\mathbf{s}^{\prime})$ and $(\mathbf{x}-\mathbf{x}_{m},s)$ be the distributions obtained after subtracting from the elements in the support of $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ and $\mathbf{(x,s)}$ the respective distribution medians. If $(\mathbf{x}^{\prime}-\mathbf{x}^{\prime}_{m},\mathbf{s}^{\prime})$ is a spread of $(\mathbf{x}-\mathbf{x}_{m},s)$, then $\hat{V}_{p}(\mathbf{x}^{\prime},\mathbf{s}^{\prime})>\hat{V}_{p}(\mathbf{x},\mathbf{s})$ for $p=A,B$.
Why do parties benefit, in equilibrium, when the conflict between voters on opposite sides of the political spectrum increases? A spread of the distribution of voters has two effects: first, if both those on the left and those on the right become more extreme, they also become less likely to vote for the party targeting the opposite voter group. Second, a greater divide between voters on the opposite ends of the political spectrum allows parties to diverge more, further securing the support of their respective electoral bases.
Redistributive Conflict and Inequality
Going back to our redistributive conflict example, proposition 3.3 tells us that politicians might benefit from increases in income inequality: a transfer of income from types in the bottom half of the income distribution to those in the top half will exacerbate the conflict over ideal taxation, spreading the distribution of tax policy bliss points. Similarly, a shrinking middle class – more probability mass in the two tails of the income distribution and less mass around the median income – would also spread the distribution of ideal tax rates, increasing both parties’ equilibrium expected utility. Finally, note that both parties benefit if the rich perceive the marginal benefit from the public good $\psi$ to be smaller than the poor. This type of correlation between income and the perceived costs and benefits from public good provision exacerbates conflict over redistribution, making the two parties better off in equilibrium.
In the previous sections, we have shown that our simple model delivers novel insights into party divergence and voter polarization. When the marginal benefit of political rents is decreasing, parties have an incentive to diverge, to secure votes: by appealing to different “core voters,” they limit competition and, as a consequence, the variability in their electoral support. Both parties benefit from alienating the extreme supporters of their competitor, and in general from making the distribution of voters more polarized. In the next section, we extend our analysis to the, perhaps more realistic, case where party platforms are multi-dimensional. We show that our main results are robust to multi-policy competition, except that in this case, the characteristics of the joint distribution of voter ideal policies matter when defining the type of increase in voter conflict that benefits parties in equilibrium, yielding additional insights on the role of multi-dimensional disagreement.
4 Multidimensional Conflict
Let us now consider $K>1$ policies, so that platforms $x_{A}$ and $x_{B}$ are vectors of $\mathbb{R}^{K}$. While we still denote a voter distribution by $(\mathbf{x},\mathbf{s})$, note that now a type $x_{i}$ is also a vector of $\mathbb{R}^{K}$. The analysis is carried out under the same assumptions of section 3, with some additional restrictions. First, we require $\nu$ to be strictly concave. Strict concavity guarantees that party best responses never fall at points where payoffs are non-differentiable, which greatly simplifies the analysis. [^17] The assumption can be relaxed and, in particular, we can allow for a jump in payoff when parties reach the majority of votes. Second, for some of the main results of the section we restrict the attention to symmetric voter distribution, which we prove to admit a pure strategy Nash equilibrium. [^18] In the more general formulation of the game, the existence of a mixed strategy equilibrium is guaranteed by the Glicksberg theorem, but the existence of a pure strategy equilibrium is not, because the utility function of a party is not quasi-concave in that party’s strategy.
Definition 3
A voter distribution $(\mathbf{x},\mathbf{s})$ is symmetric if there exists a vector $x_{m}\in\mathbb{R}^{K}$ such that for each voter group $i$ with ideal policies $x_{i}\in\mathbb{R}^{K}$ there exists a voter group $\bar{i}$ with ideal policies $x_{\bar{i}}=x_{m}-(x_{i}-x_{m})$ and such that $s_{\bar{i}}=s_{i}$.
We denote the set of all (discrete) symmetric voter distributions as $\mathcal{V}$. When symmetry is needed for the results of our proposition, we will specify that $(\mathbf{x},\mathbf{s})\in\mathcal{V}$.
A final working assumption made in this section is that voters assign the same importance to all policy dimensions so that the utility function of a voter from group $i$ takes the form $v(x,x_{i})=g(x_{i})-\alpha||x_{i}-x||^{2}$ for some $\alpha>0$ (without loss of generality we set $\alpha=1$). The assumption can be very easily relaxed to account for dimension-specific weights, capturing the relative salience of each policy. [^19] In particular, assigning weight $\alpha^{k}$ to dimension $k$ is equivalent to assuming dimensions have equal weight after replacing each voter group $i$ ideal policy $x^{k}_{i}$ with the transformed policy $\tilde{x}_{i}=(\alpha^{k})^{2}x^{k}_{i}$, and letting parties choose directly the transformed policy $\tilde{x}^{k}_{p}=(\alpha^{k})^{2}x^{k}_{p}$.
Section 4.1 discusses two equilibrium concepts that prove useful in analyzing the two-dimensional case, and presents our results on equilibrium existence as well as algorithms that can be used to find the Nash equilibria of the game. We then move to the comparative statics analysis of section 4.2, and show how changes in voters’ multidimensional conflict affect parties’ equilibrium expected payoffs.
4.1 Equilibrium Analysis
The equilibrium analysis presents some technical challenges. First, parties’ expected payoffs are not differentiable at platform profiles $(x_{A},x_{B})$ such that some voter group $i$ is indifferent between voting for party $A$ or party $B$. Second, a party’s expected payoff is not quasi-concave in that party platform – best responses are not continuous in the competitors’ platforms. Third, the game is not finite, in the sense that the strategy set is non-finite. These features complicate the search for pure strategy equilibria, as even their existence is not guaranteed by standard results. The first step of the equilibrium analysis is introducing the notion of local equilibrium of the election game.
Definition 4
A profile $(x^{l}_{A},x^{l}_{B})$ is a local equilibrium of the game if there exists a scalar $u>0$ such that $x=x^{l}_{A}$ maximizes $V_{A}(x,x^{l}_{B})$ subject to the constraint that $||x-x^{l}_{A}||<u$ and $x=x^{l}_{B}$ maximizes $V_{B}(x^{l}_{A},x)$ subject to the constraint that $||x-x^{l}_{B}||<u$. We say that $x^{l}_{A}$ is a local best response to $x^{l}_{B}$ and $x^{l}_{B}$ is a local best response to $x^{l}_{A}$.
In other words, a platform pair is a local equilibrium if neither party has an incentive to marginally adjust their platform. Our next result provides a useful characterization of the set of local equilibria of the game. Before proceeding, however, we need some additional notation.
Let $\mathcal{R}(I)$ denote the set of permutations of $I$ and for each $i\in I$ and $R\in\mathcal{R}(I)$, let $R_{i}$ denote the type in position $i$ of $R$ [^20] For instance, if $I=\{1,2,3,4\}$ and $R=(2,3,1,4)$, then $R_{3}=1$ and $R_{1}=2$. Next, define $\bar{x}:\mathcal{R}(I)\to\mathbb{R}^{K}$ and $\underaccent{\bar}{x}:\mathcal{R}(I)\to\mathbb{R}^{K}$, such that
$\bar{x}(R)=\sum_{i=1}^{N}\bar{w}_{R_{i}}x_{R_{i}}$
$\underaccent{\bar}{x}(R)=\sum_{i=1}^{N}\underaccent{\bar}{w}_{R_{i}}x_{R_{i}}$
where $\bar{w}_{R_{i}}=\nu(\bar{s}_{R_{i}})-\nu(\bar{s}_{R_{i+1}})$, $\underaccent{\bar}{w}_{R_{i}}=\nu(1-\bar{s}_{R_{i+1}})-\nu(1-\bar{s}_{R_{i}})$, and $\bar{s}_{R_{i}}=\sum_{j=i}^{N}s_{R_{i}}$. Finally for any two platforms $x,x^{\prime}\in\mathbb{R}^{k}$, let $\mathcal{r}(x,x^{\prime})\subset\mathcal{R}(I)$ contain any strict ranking of types, from types more likely to vote for $B$ to those more likely to vote for $A$, when $A$ offers $x$ and $B$ offers $x^{\prime}$. Formally, for each $R\in\mathcal{r}(x,x^{\prime})$ and $i,j\in I$, $i>j\implies\Delta_{R_{i}}(x,x^{\prime})>\Delta_{R_{j}}(x,x^{\prime})$. Note that $\Delta_{R_{i}}(x,x^{\prime})$ contains at most one element, and is empty if the considered platform profile makes two voter groups ex-ante indifferent.
Equipped with this notation we are ready to state the next result: if platforms $\bar{x}(R)$ and $\underaccent{\bar}{x}(R)$ preserve the ordering $R$ if and only if they are locally optimal, as described in the next lemma.
Lemma 4.1
Let $\mathcal{L}^{p}$ and $\mathcal{L}$ be the set of local best responses and local equilibria of the game, respectively, and let
$\mathcal{L}^{p}_{\mathcal{R}}=\left\{x\in\mathbb{R}^{K}:\exists R\in\mathcal{R}(I),x\in\{\bar{x}(R),\underaccent{\bar}{x}(R)\}\right\}$
=$\mathcal{L}_{\mathcal{R}}=\left\{(x_{A},x_{B})\in\mathbb{R}^{2K}:\exists R\in\mathcal{r}(x_{A},x_{B}),(x_{A},x_{B})=(\bar{x}(R),\underaccent{\bar}{x}(R))\right\}.$
It has
$\mathcal{L}^{p}\subseteq\mathcal{L}_{\mathcal{R}}^{p}$
$\mathcal{L}=\mathcal{L}_{\mathcal{R}}.$
Moreover, for each local equilibrium $(x_{A},x_{B})\in\mathcal{L}$, each party’s expected payoff $V_{p}(x_{A},x_{B})$ is proportional to $||x_{A}-x_{B}||^{2}$ up to a constant term.
An immediate implication of the lemma is that, to find the set of local equilibria, it is sufficient to look for the rankings $R$ of voter types, from core supporters of $B$ to core supporters of $A$, which are preserved after parties propose the policies that are consistent with such voter segmentation. Moreover, the payoff achieved by parties in equilibrium increases with the distance between party platforms: as for proposition 3.1, the intuition is that the more different the parties are, in equilibrium, from each other, the more likely they will be to win the support of at least one end of the political spectrum. From an instrumental perspective, the lemma is also important because $\mathcal{L}^{p}\subseteq\mathcal{L}^{p}_{\mathcal{R}}$ reveals that, for each player, only a finite number of platforms can ever be local best responses. [^21] In fact, each ordering $R\in\mathcal{R}(I)$ admits unique $\bar{x}(R)$ and $\underaccent{\bar}{x}(R)$, and the number of orderings (permutations) is finite, $|\mathcal{R}(I)|<\infty$, since $|I|<\infty$. We can therefore constrain players to choose pure strategies in the finite set $\mathcal{L}^{p}_{\mathcal{R}}$, obtaining a game with the same local equilibria as the original game.
The notion of local equilibrium is meaningful in contexts where large policy changes are unreasonable – for instance because they are more costly to design, not credible or politically not viable. When this is not the case, local equilibria are less interesting than Nash equilibria, as they do not rule out the possibility that at least one party could benefit from changing policy platforms dramatically, in a way that causes realignments of voters.
The next result plays a key role in establishing the characteristics of pure strategy Nash equilibria in the case of symmetric voter distributions.
Lemma 4.2
Let $(\mathbf{x},\mathbf{s})\in\mathcal{V}$, and let $x_{0}\in\mathcal{L}$, $x_{0}=(x_{A0},x_{B0})$. For $n\in\mathbb{N}$ let $x_{n}$ be the profile obtained in the $n^{th}$ step of the best response dynamics initiated at the local equilibrium $x_{0}$. Then $||x_{An}-x_{Bn}||^{2}\geq||x_{A0}-x_{B0}||^{2}$. The inequality holds strictly after a finite number of iterations if $x_{0}$ is not a Nash equilibrium.
If a party has an incentive to deviate from a local equilibrium, letting parties sequentially best respond to each other after the deviation will keep them further apart than before the deviation. In other words, if the sequence of policy adjustments converges, the resulting platform will make both parties better off relative to the one that has been abandoned. This intuition allows us to find a characterization for a set of focal Nash equilibria, those that maximize the expected value of the election game for each politician.
Let $\mathcal{E}$ be the set of pure strategy equilibria of the game and define the set of party preferred equilibria as follows,
:-$\mathcal{E}^{\star}\coloneq\arg\max_{(x_{A},x_{B})\in\mathcal{E}}V_{p}(x_{A},x_{B}).$
Since each party gets the same expected payoff in equilibrium, $\mathcal{E}^{\star}$ does not depend on $p$. The following proposition tells us that a pure strategy Nash equilibrium of the game exists, and provides a simple characterization of $\mathcal{E}^{\star}$
Proposition 4.1
Let $(\mathbf{x},\mathbf{s})\in\mathcal{V}$. The set of party-preferred pure strategy Nash equilibria is non-empty and is the set of local equilibria where parties differ the most in terms of their proposed platforms.
$\mathcal{E}^{\star}=\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}\neq\emptyset.$
In the appendix, we show that, under our assumptions, the game admits a potential function. The rest of the proposition ties together this observation with the main results of this section. First, lemma 4.1 implies that the strategy space is essentially finite so that the best response dynamics converge to a pure strategy equilibrium (the potential function can only increase a finite number of times). Second, the set of party-preferred Nash equilibria is the set of party-preferred local equilibria, and Nash equilibrium platforms must be as different as they can get when local equilibrium platforms are considered. [^22] To see why, note that if a profitable deviation from such profile existed, by lemma 4.2 we should be able to find a local equilibrium where parties’ platforms are even more different, by simply applying best response dynamics. So the local equilibrium that maximizes the distance between party platforms – and therefore party payoffs – is also the party-preferred Nash equilibrium.
To summarize, in this section, we proved the existence of a pure strategy Nash equilibrium for the game with $K$ dimensions and a symmetric voter distribution. Equilibria can be ranked in terms of party equilibrium expected payoff (the ranking is the same regardless of which party is considered). Given this common interest on the policy supply side, the focal Nash equilibria, in our analysis, are the ones that grant the highest expected payoffs to parties. Not surprisingly, in these equilibria, parties are maximally differentiated, confirming the intuition of the previous sections: platform divergence allows parties to better target their most likely voters, securing at least some office rents.
4.1.1 Algorithmic Solutions
The task of finding the local equilibria is simplified by our result of lemma 4.1. In fact, the following procedure can be used to construct $\mathcal{L}$ in a finite number of steps. First, let $\mathcal{L}_{0}=\emptyset$, construct $\mathcal{R}(I)$ as the set of permutations of $I$, and let $R^{1},R^{2},...,R^{|\mathcal{R}(I)|}$ denote the elements of $\mathcal{R}(I)$. Second, for each $r=1,...,|\mathcal{R}(I)|$, perform the following steps: (i) compute $\bar{x}(R^{r})$ and $\underaccent{\bar}{x}(R^{r})$; (ii) compute $\mathcal{r}(\bar{x}(R^{r}),\underaccent{\bar}{x}(R^{r}))$; (iii) Let $\mathcal{L}_{r}=(\{R_{r}\}\cap\mathcal{r}(\bar{x}(R^{r}),\underaccent{\bar}{x}(R^{r})))\cup\mathcal{L}_{r-1}$. By lemma 4.1, $\mathcal{L}_{|\mathcal{R}(I)|}=\mathcal{L}$. Furthermore, if $(\mathbf{x},\mathbf{s})$ if symmetric, proposition 4.1 tells us that we can easily find the party preferred Nash equilibria by simply looking for the elements of $\mathcal{L}$ that maximize party expected utility $V_{p}$.
(a) Asymmetric
(b) Symmetric
Figure 1: Local and Party-Preferred Equilibria
[…]
Notes. Panel (a) plots a non-symmetric bidimensional voter distribution and the associated local equilibria where party payoffs. Panel (b) plots a symmetric bidimensional voter distribution and the associated local equilibria. In both panels: (i) ”x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) two circles of the same color represent a pair of local equilibrium platforms; (iii) we assume that party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$.
Figure 1 panel (a) reports a voter distribution over a bidimensional policy space, as well as the set of local equilibria, found applying the algorithm outlined in the previous paragraph. Lemma 4.1 tells us that among the existing local equilibria, the one identified by the two green dots is associated to the highest party expected payoff since the platforms proposed by the two politicians differ the most. In panel (b) we report the voter types and local equilibria for a symmetric voter distribution. In this case, the local equilibrium in blue is also the Nash equilibrium where parties expect to benefit the most, as per the results of proposition 4.1.
In section 3.1 we found that, with one policy dimension, politicians benefit from pushing voters to the left of the median voter further to the left, and those to the right further to the right. In the next section, we show that a similar intuition holds when multiple policy dimensions are at stake: the political supply benefits when the population grows increasingly divided into two factions of voters, disagreeing on multiple policy issues simultaneously.
4.2 Divide and Diverge
As mentioned in the introduction, one of the most striking patterns of today’s political conflict is the existence of highly polarized factions of voters: views of the American electorate have become more correlated across policy issues (e.g., Gentzkow,, 2016; Bonomi et al.,, 2021; Pew Research Center,, 2023). This section provides new insights on why these types of changes might benefit the political supply side, shedding light on the connection between multidimensional disagreement and the (unidimensional) voter polarization discussed in the previous sections. As in section 3.1, we will express the objects of interest as functions of the voter distribution considered, so that, for instance, $\mathcal{L}(\mathbf{x},\mathbf{s})$ and $\mathcal{E^{\star}}(\mathbf{x},\mathbf{s})$ denotes the sets of local equilibria and party preferred equilibria associated to voter distribution $(\mathbf{x},\mathbf{s})$, respectively.
We start from a result that applies to our local analysis: for each local equilibrium, we identify the marginal change of opinions of a voter group that benefits parties the most. Not surprisingly, parties benefit when voters at opposite ends of the political spectrum become even more different. Before illustrating the result in the next proposition, we need to define a few more concepts. For any local equilibrium $x\in\mathcal{L}$, with associated voter ranking $R\in\mathcal{r}(x)$, define $m_{R}\in\mathbb{R}$ to be such that $\sum_{i<m_{R}}s_{R_{i}}<\frac{1}{2}$ and $\sum_{i>m_{R}}s_{R_{i}}<\frac{1}{2}$. For $i>m_{R}$, we say that group $R_{i}$ is a relative supporter of party $A$, while for $i<m_{R}$ then we say that $R_{i}$ is a relative supporter of party $B$. When $x_{p},x_{-p}\in\mathbb{R}^{K}$ are played by the two parties, we say that $x_{p}-x_{-p}$ is the characteristic direction of party $p$, and, similarly, $x_{p}^{k}-x_{-p}^{k}$ is the characteristic direction of party $p$ for policy $k$. Finally, let $\hat{V}^{R}_{p}(\mathbf{x},\mathbf{s})$ denote party $p$’s expected payoff when the voter distribution is $(\mathbf{x},\mathbf{s})$ and the local equilibrium inducing $R$ is played.
Proposition 4.2
Fix a voter distribution $(\mathbf{x},\mathbf{s})$ and a local equilibrium ordering $R$. The following holds about the effect of marginal changes in voter opinions on parties’ equilibrium payoffs.
  1. (i)
    Consider any policy $k=1,...,K$. Both parties benefit if the ideal policy $k$ of a relative supporter of party $p=A,B$ changes marginally in $p$’s characteristic direction. Formally, for each $i\in I$,
    $\frac{\partial\hat{V}^{R}_{p}}{\partial x_{R_{i}}^{k}}(\bar{x}^{k}(R)-\underaccent{\bar}{x}^{k}(R))>(<)\text{ }0\text{ if }i>(<)\text{ }m_{R}$
  2. (ii)
    Parties benefit the most if the marginal change of opinions of the supporters of party $p=A,B$ is in the direction characteristic of party $p$.
We illustrate these results using the example of a bi-dimensional policy competition over the size of government and abortion policy. In a local equilibrium, party $A$ appeals to conservative right supporters, both against redistribution and against abortion, while party $B$ targets the opposite ends of the political spectrum, offering policies that are more progressive on both issues than its political opponent.
A consequence of the first part of proposition 4.2 is that both political parties benefit, in the local equilibrium, if any group of party $B$ supporters becomes even more in favor of abortion or redistribution, with a (marginal) increase of disagreement with those who support party $A$. The second part of the proposition suggests that parties benefit the most when the two groups of supporters become marginally more different in both policy dimensions simultaneously, with an increase in disagreement on each policy issue that is proportional to how much parties are already differentiated on such issue.
The intuition is analogous to the one for unidimensional competition (section 3). When the progressive strata of the electorate become even more distant from their conservative counterpart, two things happen: first, the relatively more progressive party benefits, as its core supporters are less likely to switch to the other party (a switch will only occur for more extreme popularity shocks). Second, the progressive party – which over-weights the views of its core supporters – offers a platform that is more progressive than before. This platform change further alienates conservative voters, who are less likely to swing away from the most conservative party, which therefore also benefits from the increase in conflict.
Figure 2: Directions of Opinion Change for Fastest Payoff Improvement
[…]
Notes. The figure plots bidimensional voter distribution and the associated local equilibria. In the plot, ”x” markers represent voter types’ ideal policies, and each voter group is equally sized; two circles of the same color represent a pair of local equilibrium platforms; we assume that the party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$. For each group of party supporters, red arrows identify the direction of opinion change of party supporters that yields the fastest party payoff improvement in local equilibrium 1. Similarly, for each group of party supporters, blue arrows identify the direction of opinion change of party supporters that yields the fastest party payoff improvement in local equilibrium 2.
Figure 2 reports the direction in which voter opinions should change for parties’ equilibrium payoffs to increase the fastest, for the two local equilibria of figure 1, panel (a). The red arrows represent this direction for the best local equilibrium (identified by the green circles), while the blue arrows identify the direction for the other local equilibrium (identified by the white circles). In both cases, the change makes supporters of party $A$ on average more different from supporters of party $B$.
The results of proposition 4.2 hold for an arbitrary distribution of voter types, as long as the policy opinions changes considered are marginal and the equilibrium payoffs compared before and after the change are those achieved in two local equilibria that induce the same voter ranking $R$. However, after larger distributional changes, some local equilibria might cease to exist. Moreover, as already mentioned, the focus on local equilibria is less justifiable when large policy changes are feasible. The next proposition extends the intuition of proposition 4.2 to non-local changes in the voter distribution and party-preferred Nash equilibria, under the constraint that the voter distribution considered is symmetric. Before stating the result, it helps to define the concept of $d$-spread of a voter distribution.
Definition 5
Let $d\in\mathbb{R}^{K}$, where $K$ is the number of policy dimensions. Distribution $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ is a $d$-spread of distribution $\mathbf{(x,s)}$ if $(\mathbf{x}^{\prime T}d,\mathbf{s}^{\prime})$ is a spread of $(\mathbf{x}^{T}d,\mathbf{s})$.
In other words, a distribution of voter ideal policies is a $d$-spread of another if, after the elements of their support are projected on $d$, the former’s distribution of projection coefficients is a spread of the latter.
[…]
Proposition 4.3
Both parties benefit from changes in voters’ views that exacerbate conflict along the party-preferred equilibrium direction. Formally, consider two voter distributions $(\mathbf{x},\mathbf{s})$,$(\mathbf{x}^{\prime},\mathbf{s}^{\prime})\in\mathcal{V}$, and let $\tilde{x}\in\mathcal{E}^{\star}(x,s)$. If $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ if a $d$-spread of $(\mathbf{x},\mathbf{s})$, for $d=\tilde{x}_{A}-\tilde{x}_{B}$, then $\hat{V}_{p}(\mathbf{x}^{\prime},\mathbf{s}^{\prime})>\hat{V}_{p}(\mathbf{x},\mathbf{s})$ for $p=A,B$.
The proposition has the following interpretation. Fix an initial voter distribution and any associated party-preferred Nash equilibrium. Any change in the voter distribution that makes at least some supporters of party $p$ relatively more in favor of such party than before increases the expected payoff achieved by both parties in the new party-preferred Nash equilibrium.
Note that the same insights as for marginal changes in voter opinions hold also when larger changes are considered. First, if all voters that support a more right-wing party shift further to the right in terms of their views about redistribution, parties will be better off, because, in the new equilibrium, they can offer fiscal agendas that are more differentiated (i.e. they can diverge more on the economic dimension). The effect of this change in the marginal distribution is illustrated in panels (a) and (b) of figure 3. Second, a larger payoff increase is obtained if voter views change in the direction that is characteristic of their preferred party, given the initial party preferred equilibrium $\tilde{x}$: the growing disagreement on multiple policy dimensions at a time makes voters more unlikely to switch faction.
Besides extending our previous intuitions beyond marginal changes, proposition 4.3 yields a new, interesting, insight. It tells us that even an increase in the cohesiveness of the two factions of voters – i.e., an opinion change that leaves unaffected the marginal distribution of views on abortion and redistribution but increases their degree of correlation – will on average benefit parties. Consider the distribution $(\textbf{x}^{\prime},\textbf{s})$, obtained from $(\mathbf{x},\mathbf{s})$ as follows: for each $i\in I$ and $k\leq K$, let $x_{i}^{\prime k}=x_{i}$ if $\Delta_{i}(\tilde{x})-\Delta_{\bar{i}}(x)>0$ and $(x^{k}_{i}-x^{k}_{\bar{i}})(\tilde{x}^{k}_{A}-\tilde{x}^{k}_{B})\geq 0$ and $x_{i}^{\prime k}=x^{k}_{\bar{i}}$ otherwise. It is immediate to see that $(\textbf{x}^{\prime},\textbf{s})$ and $(\mathbf{x},\mathbf{s})$ have the same marginal distributions, but it can be shown the whenever the two joint distributions are different, the former is a $\tilde{x}$-spread of the latter.
Remark 4
Increases in the ideological cohesiveness of opposite voter factions increase parties’ preferred equilibrium payoffs.
The opposite groups of party supporters disagree more than before because some of the party $B$ supporters that originally had pro-party $A$ views on policy issue $k$ have changed their mind on that policy in the direction that is characteristic of party $B$, and a similar realignment as occurred for the supporters of party $A$. While polarization over single policy issues is left unaffected by the opinion change, the two factions disagree on more topics.
(a) Marginal Distribution Change (MDC)
(b) MDC: Party Readjustment
(c) Ideological Cohesion Change (ICC)
(d) ICC: Party Readjustment
Figure 3: Changes in Voter Conflict
[…]
Notes. In both panels: (i) ”x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) green circles represent the party-preferred equilibrium platforms; (iii) we assume that the party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$. Panel (b) plots the voter distribution obtained starting from panel (a) and moving selected voter types in the direction identified by the red arrows in panel (a). The white circles of panel (b) represent the party-preferred equilibrium profile of panel (a), before the distributional change. Panel (d) plots the voter distribution obtained starting from panel (c) and moving selected voter types in the direction identified by the red arrows of panel (c). The white circles of panel (d) represent the party-preferred equilibrium profile of panel (c), before the distributional change.
Panels (c) and (d) of figure 3 illustrate the argument in a bi-dimensional space. For interpretation purposes, let the horizontal axis represent the redistribution policy dimension (markers on the right represent more right-wing policies), and let the vertical axis represent abortion policy (higher markers representing more liberal policy positions).
In panel (c), the correlation between voters’ views on abortion and redistribution is almost null, and Nash equilibrium policies differ mostly on abortion. The culturally conservative party proposes moderately less redistribution than the culturally progressive: the two parties over-weight the opinions of their extreme supporters, who are more economically conservative and progressive, respectively. In this configuration, parties’ economic agendas collide with the demands of some of their voters: indeed there are some anti-abortion voters are on the left of the economic policy spectrum, while some of the more progressive voters belong to the economic right.
The voter distribution of panel (d) is obtained from panel (c) by switching the economic views of the two groups of “dissonant” supporters, with an opinion change in the directions highlighted by the red arrows. In panel (d), voters appear more cohesive: the right-wing party supporters are all conservative on both policy dimensions, while the opposite is true for the supporters of the left-wing party. The increased economic conflict between the two factions allows politicians to differentiate more their fiscal policy proposals, further building their voter loyalty and gaining in expected payoff.
Main Takeaways on Voter Conflict
In section 3.1 we showed that, with one policy dimension, parties benefit if voters’ views become more polarized (in a “spread” sense). Extending the analysis to a multidimensional setting, we have obtained a more general insight: what benefits parties in equilibrium is an electorate fractured into two factions of voters, very different in their policy demands. While a spread of marginal distributions of policy views is no longer guaranteed to advantage the political supply side if multiple policy issues are at stake, an increase in the overall disagreement between these two factions, with voters from one group becoming more ideologically distant from those in the other group, does result in a gain for both political parties. This increased divide in the electorate can come in the form of greater “ideological cohesion” of the two factions, with marginal distributions of policy views remaining unaffected but views becoming more correlated together.
The above results, taken together, have a natural implication: parties could have an incentive to divide voters. In the next section, we interpret themes on party identity, zero-sum thinking, and common interest vs conflict of interest policies in light of these incentives.
5 Applications
In this section, we briefly discuss applications of our theory to zero-sum policies, parties’ information provision incentives, and party identity.
5.1 Common Interest vs Conflict of Interest Issues
A recent stream of literature in political science and economics has stressed how politicians tend to focus on divisive, zero-sum, issues over common interest issues, even if the latter entail policies that could benefit a larger share of voters (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023). This phenomenon is potentially amplified by the fact that, on the one hand, voters seem to be prone to social and political “zero-sum thinking,” the perception that gains from one social group must necessarily come at the expense of another (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), and often lack the information to assess the “common value” nature of policies (Matsusaka and Kendall,, 2022), while politicians’ rhetoric often paints politics in a zero-sum fashion. Our simple model can provide a new perspective on why politicians might have an incentive to focus on more divisive issues, stress the zero-sum component of policy-making, and keep voters informed on zero-sum issues and uninformed about common interest policies.
We illustrate these points through a parsimonious version of our model, where there are only two voters. One voter is religious ($R$), the other is secular ($S$), and there is one policy domain, say government spending aimed at tackling climate change. Consider now two scenarios. In the common interest scenario, tackling climate change is interpreted as an issue where interests are aligned, so that voters have the same bliss point, $x_{R}=x_{S}$. In the conflict of interest scenario, the religious voter interprets any policy aimed at addressing climate change as a negation of God’s control over the climate, [^23] As documented by Pew Research Center, (2022), such an interpretation is common among some religious voters. so that she is less supportive of such a spending item relative to the secular voter, $x_{R}<x_{S}$. An immediate consequence of the analysis of the previous sections is that, in equilibrium, politicians are better of in the conflict of interest scenario. If $x_{R}=x_{S}$, then $\bar{x}=\underaccent{\bar}{x}$, that is, the two candidates are forced to converge on the ideal policy that benefits both groups, obtaining an equilibrium expected payoff $\frac{\nu(1)+\nu(0)}{2}$; if instead $x_{R}<x_{S}$, in equilibrium the two candidates target opposite voters, and differentiate in terms of the proposed environmental spending, with an equilibrium expected payoff of $\frac{\nu(1)+\nu(0)}{2}+\frac{(\bar{x}-\underaccent{\bar}{x})^{2}}{2\phi}>\frac{\nu(1)+\nu(0)}{2}$. In general, we find that candidates benefit when voters perceive the policies at stake as advantaging some groups over others: in that scenario, politicians can cater to different interests in the electorate rather than competing on the same votes.
Remark 5
Candidates benefit if voters perceive a policy issue as defining a conflict of interest in society, even when in reality the same policy could benefit everyone.
Following the same type of argument, our model suggests that the political supply side should have an incentive to downplay the importance of common interest issues relative to issues where different groups have different policy goals. We illustrate this point with our two voters, $S$ and $R$, by introducing a second policy instrument $q$, a healthcare reform. Assume that the utility of voter group $i$ takes the form
$v(x,q,x_{i},q_{i})=-\alpha(x_{i}-x)^{2}-(1-\alpha)(q_{i}-q)^{2},$ (2)
where $\alpha$ captures the relative payoff importance of policy $x$. Assume further that $x_{R}\neq x_{S}$ and $q_{R}=q_{S}=q^{\star}$ so that the former policy is a conflict of interest policy and the latter is a common interest policy. By our previous results, in equilibrium candidates offer the same platform $q^{\star}$ for the common interest issue, but offer different policies, $\bar{x}$ and $\underaccent{\bar}{x}$, for the controversial issue. Their equilibrium utility is $\frac{\nu(1)+\nu(0)}{2}+\alpha\frac{(\bar{x}-\underaccent{\bar}{x})^{2}}{2\phi}$. Hence, candidates’ payoffs are increasing in $\alpha$, the relative payoff importance of the conflict of interest issue: politicians have an incentive to render such an issue more salient, increasing $\alpha$, making the zero-sum component politics more salient. [^24] Note that the same would be true if $x$ and $q$ were zero-sum and common interest components of the same policy issue.
Remark 6
Candidates have an incentive to render conflict of interest policy issues more salient than common interest ones.
Finally, note that the same incentive to divide voters predicts an under-provision of information on common interest policies and an over-provision of information on zero-sum policies. Let voters have the utility in 2 and assume that voters and politicians are uncertain about $q^{\star}$ and have the common prior that $q^{\star}=1$ with probability $\pi_{q}\in(0,1)$, and $q^{\star}=0$ with probability $1-\pi_{q}$. Similarly, assume that voters and politicians believe that with probability $\pi_{x}\in(0,1)$ it has $x_{R}=1$ and $x_{S}=0$, while with probability $1-\pi_{x}$, $x_{R}=0$ and $x_{S}=1$, [^25] Voters’ uncertainty about their ideal policy stems from uncertainty about an underlying state. and these events are independent of $q^{\star}$.
Consider now the policy platforms $q^{o}$ and $x^{o}$ that maximize voters’ utilitarian welfare. Two observations are due. First, $q^{o}=q^{\star}$, that is, as far as the common interest issue is concerned, the same policy can make all voters maximally well off. Second, $x^{o}=\frac{1}{2}$, since, by the concavity of voters’ payoffs the $x$-policy that maximizes their aggregate welfare lies between $x_{R}$ and $x_{S}$.
Next, let us examine how information influences voter welfare and politicians’ payoffs, in the equilibrium of the party competition game. To keep things simple, assume that, for each policy separately, either politician can pay a small cost to reveal the state before the competition game starts. Uncertainty is easily accommodated in our framework: after updating their beliefs based on any information that was transmitted, voter $i$ casts a vote in favor of politician $A$ if and only if $\mathbb{E}\left[\Delta_{i}\right]\leq\epsilon$.
Consider policy $q$ first. If any of the two politicians pay the small cost, voters learn $q^{\star}$, and in equilibrium $q_{A}=q_{B}=q^{\star}$. If instead no information on $q^{\star}$ is acquired, in equilibrium $q_{A}=q_{B}=\pi_{q}$. For any policy pair $\bar{x}$ and $\underaccent{\bar}{x}$ implemented in equilibrium, the (expected) utilitarian voter welfare is larger in the equilibrium where $q^{\star}$ is learned, with an expected utility gain of $(1-\alpha)\pi_{q}(1-\pi_{q})$ per voter relative to the case where no information is obtained, due to the uncertainty reduction. However, parties’ equilibrium expected utility does not increase when $q^{\star}$ is learned, because obtaining information does not change their degree of differentiation on policy $q$ (i.e., they converge regardless of the amount of information obtained). Parties have no incentive to acquire and provide information to voters about common interest issues, even if such information improves voter welfare.
Remark 7
Candidates have no incentive to keep the electorate informed about common interest issues, even when such information is welfare-improving.
Now consider policy $x$. Without loss of generality, we let party $A$ be the one targeting the $R$ voter in equilibrium, so that it has
$\bar{x}=\left[\nu\left(\frac{1}{2}\right)-\nu\left(0\right)\right]\hat{\pi}_{x}+\left[\nu\left(1\right)-\nu\left(\frac{1}{2}\right)\right](1-\hat{\pi}_{x}),$
$\underaccent{\bar}{x}=\left[\nu\left(\frac{1}{2}\right)-\nu\left(0\right)\right](1-\hat{\pi}_{x})+\left[\nu\left(1\right)-\nu\left(\frac{1}{2}\right)\right]\hat{\pi}_{x},$
$\bar{x}-\underaccent{\bar}{x}=\gamma(2\hat{\pi}_{x}-1),$
where $\hat{\pi}_{x}$ is the probability that $x_{R}=1$ according to the common belief held by the voter when candidates select their platforms and $\gamma=2\nu\left(\frac{1}{2}\right)-\left(\nu(1)+\nu\left(0\right)\right)>0$. First, note that if politicians pay the cost to inform voters about the $x$-relevant state, then $\hat{\pi}_{x}\in\{0,1\}$, while if they do not pay such cost it has $\hat{\pi}_{x}=\pi_{x}\in(0,1)$. Second note that, the further $\hat{\pi}_{x}$ is from $\frac{1}{2}$, the further both $\bar{x}$ and $\underaccent{\bar}{x}$ will be from the welfare-maximizing policy $x^{o}=\frac{1}{2}$. When the state is revealed, voters $R$ and $S$ become more extreme in their policy demands (equal to $\hat{\pi}_{x}$ and $1-\hat{\pi}_{x}$ respectively), parties offer policies that are further apart from the welfare optimum, and the utilitarian welfare decreases relative to the case of no information provision. [^26] While we have not specified the details of the functioning of the political system after the election occurs, the implication on welfare holds provided that the implemented platform is either $\bar{x}$ or $\underaccent{\bar}{x}$, or more in general, as long as it is different from $\frac{1}{2}$ with positive probability. At the same time, in this case providing information benefits candidates, as it allows them to differentiate more from each other in equilibrium: when the state is revealed, in equilibrium it has $(\bar{x}-\underaccent{\bar}{x})^{2}=\gamma^{2}$, which is greater than $\gamma^{2}(2\pi_{x}-1)^{2}$, the squared difference in equilibrium platforms attained when the two politicians do not acquire and disclose the policy-relevant information. This finding is in stark contrast with our result for the common interest issue: parties have the incentive to provide voters with more information about divisive issues, even when doing so is welfare detrimental.
Remark 8
Candidates have the incentive to keep the electorate informed about conflict of interest issues, even when such information is welfare detrimental.
5.2 Party Identity
What role is played by party identity in this picture? As with other forms of social identity, party identity contributes to defining voters’ perceptions of their selves and others based on the party that they support (Greene,, 1999). The political implications of social identities based on demographic and ascriptive characteristics (e.g., race, religiosity, income) have been a recent object of study within the economic discipline (see, for instance, Bonomi et al.,, 2021; Gennaioli and Tabellini,, 2023; Shayo,, 2009), which highlights how identity can influence voters’ beliefs and behavior. In particular, Bonomi et al., (2021) proposes a model where the society is divided in two-group partitions. When voters identify with one of the corresponding groups, their beliefs are slanted toward positions that are more representative of that group relative to the outgroup (the complement societal group). Building on this model, when voter $i$ identifies with group $G$, the following reduced form equation can be easily shown to capture the implications of identity for her ideal policy $\tilde{x}_{i}$ when rational beliefs about ideal policies are normal distributions:
$\tilde{x}_{i}=x_{i}+\theta(x_{G}-x_{\bar{G}}),$ (3)
where $x_{i}$ is the rational ideal policy of voter type $i$, $x_{G}$ is the ideal policy that is typical of identity group $G$, $x_{\bar{G}}$ is the ideal policy that is typical of $i$’s identity outgroup, and $\theta>0$ captures the strength of the effects of identity on voters’ political views.
This setup can be easily adapted to provide a stylized explanation of why parties could benefit from favoring a “party-based identity,” that pushes opposite partisans away from each other. Let voters be distributed according to $(\textbf{x},\textbf{s})\in\mathcal{V}$ and parties play their preferred equilibrium $\hat{x}\in\mathcal{E}^{\star}$. Under this equilibrium, it seems sensible to assume that voters in group $i$ identify with party $A$ (i.e., $G=A$) if $\Delta_{i}(\hat{x})>0$ and with party $B$ if $\Delta_{i}(\hat{x})<0$ (and no party identity prevails if $\Delta_{i}(\hat{x})=0$).
[…]
The results of section 4.2 suggest that activating respondents’ party identity is guaranteed to increase politicians’ expected payoffs if $x_{G}=\hat{x}_{G}$ for $G=A,B$, so that typical party positions are pinned down by the respective initial equilibrium platforms. Note that in this case party identity is also efficient, in that it distorts voter opinions in the direction of the fastest party payoff increase, $\hat{x}_{G}-\hat{x}_{\bar{G}}$. In both cases, party identity exacerbates the divide between the already disagreeing factions voters, making them more loyal to their preferred party, and allowing rival politicians to better differentiate their political proposals and avoid large swings of supporters.
Remark 9
Party identity increases disagreement between opposite-party supporters in a way that benefits candidates in equilibrium.
These findings are complementary to those of Gennaioli and Tabellini, (2023), who show that, whenever opposite social identities (e.g., conservative vs progressive, or upper class vs lower class, etc.) are prevalent among the (exogenous) core voters of the competing parties, each party benefits from emphasizing one of the social identities of their core voters, provided that such practice does not backfire, equally alienating the voters of the opposite party. Here, we abstract away from cases where multiple identities are available and show instead that investing in the formation of party identities is particularly beneficial precisely because it can fracture the electorate, pushing both sides further apart from each other. As discussed in section 3.2, in our model each party benefits from propaganda even when it strengthens the party identity of its opponents, making them more extreme. Our findings on party identity (and their immediate extensions to other social identities) could help explain why parties leverage identity politics even if – as argued by some scholars (e.g., Fukuyama,, 2018; Chua,, 2019; Norris and Inglehart,, 2019) – it might come at the cost of generating a backlash among non-targeted groups, further fracturing the electorate.
5.2.1 A Dynamic Application: Self-reinforcing Polarization
To conclude the section, we briefly illustrate how party identity can generate a feedback loop between party and voter polarization, whereby a growingly divided electorate causes parties to diverge, and platforms’ differences amplify voter conflict through party identity.
For simplicity, we focus on the case of two voters, $L$ and $R$, and one policy instrument $x$, with $x_{R}<x_{L}$ representing voters’ rational bliss points. Elections are repeated in each $t=0,1,2,...$, an infinite horizon, and $x^{t}_{A}$ and $x_{B}^{t}$ denote the platforms proposed by party $A$ and $B$, respectively, at time $t$. In each period, each party can invest $c$ in a campaign highlighting partisan differences, making party identity salient.
At $t=1,2,...$, the bliss point of a voter $i=L,R$ identifying with party $G=A,B$ in that period is $\tilde{x}^{t}_{i}=x_{i}+\theta^{t}(x^{t-1}_{G}-x^{t-1}_{\bar{G}})$, where $G=A$ and $\bar{G}=B$ if and only if $\Delta_{i}(x^{t-1}_{G},x^{t-1}_{\bar{G}})>0$ as already discussed. If both parties invest in the identity campaign it has $\theta^{t}=\theta_{H}>0$, if only one does it has $\theta^{t}=\theta_{L}$ with $\theta_{H}>\theta_{L}>0$, and otherwise $\theta^{t}=0$. In period $t$, after making the campaign investment decision, the two candidates compete in the election game of section 3, for given voters’ bliss points $\tilde{x}^{t}_{L}$ and $\tilde{x}^{t}_{R}$. For simplicity, but with little loss, we assume that candidate campaign investment and platform choices are myopic, and that $\tilde{x}^{0}_{i}=x_{i}$ for $i=L,R$.
It is easy to see that when
$(x_{R}-x_{L})^{2}>\frac{\phi c}{2\gamma^{3}}\max\left\{\frac{1}{(\theta_{H}-\theta_{L})[1+\gamma(\theta_{H}+\theta_{L})]},\frac{1}{\theta_{L}[1+\gamma\theta_{L}]}\right\}$ (4)
for $\gamma=2\nu(\frac{1}{2})-\nu(1)-\nu(0)$, then on the equilibrium path both parties will pay the identity investment cost in every period $t>0$. As highlighted in the previous section, party identity allows parties to divide the electorate and offer different platforms in equilibrium. At the same time, when each voter identifies with their favorite party, increasing party differences feed back into voters’ opinions, amplifying voter polarization. Indeed, if 4 holds, it has
$|x^{t}_{A}-x^{t}_{B}|=|x^{0}_{A}-x^{0}_{A}|\sum_{s=0}^{t}\left(2\theta_{H}\gamma\right)^{s}$
$|\tilde{x}^{t}_{L}-\tilde{x}^{t}_{R}|=|x^{0}_{L}-x^{0}_{R}|\sum_{s=0}^{t}\left(2\theta_{H}\gamma\right)^{s}$
so that both party and voter polarization increase over time.
Remark 10
The incentive to divide voters through party identity can create a feedback loop between party and voter polarization.
6 Some Words on Welfare
[…]
One way to reduce such loss is reducing the spread of the voter distribution, so that it becomes more concentrated around its mean, and party platforms closer to the utilitarian optimum. However, as we have shown in the previous sections, parties’ incentives go in the opposite direction: parties want to increase voter disagreement. Another interesting way to reduce party divergence, bringing platforms closer to $x^{o}$, is by increasing the majority premium, that is, the political power (and rents) of the party winning the majority of votes. To see why this is the case, note that the weights $\omega_{i}$ and $w_{i}$ that parties attribute in equilibrium to the policy demand $x_{i}$ of voter $i$ are increasing in the rents increase that they obtain if they win $i$’s electoral support. In particular, let
$R_{m}=\lim_{s\rightarrow\frac{1}{2}^{+}}R(s)-\lim_{s\rightarrow\frac{1}{2}^{-}}R(s)$
be the majority premium. The weights $w_{m}$ and $\omega_{m}$ assigned in equilibrium to the median voter policy view $x_{m}$ are increasing in $R_{m}$, and, indeed
$\lim_{R_{m}\rightarrow\bar{R}^{-}}\bar{w}_{m}=\lim_{R_{m}\rightarrow\bar{R}^{-}}\underaccent{\bar}{w}_{m}=1$
so that, when $R\to\bar{R}$, $\bar{x}\to x^{o}$ and $\underaccent{\bar}{x}\to x^{o}$. This result is captured by the following remark.
Remark 11
Increasing the majority premium makes the equilibrium platforms converge to the median voter policy views. This results in higher utilitarian welfare if the median voter policy views are close enough to the welfare optimum.
Naturally, considerations other than political polarization come into play when assessing the benefits of a lower vs higher majority premium. Certainly, the absence of checks and balances might result in a tyranny of the majority (Mill,, 1859). The main takeaway of this section, however, is that the way votes map into political power allocation might play an important role in driving political polarization, with greater disproportionality between votes and power being associated with lower levels of platform divergence, as found, for instance, by Matakos et al., (2016).
7 Conclusion
We have proposed a novel explanation for party divergence that results in a political-supply incentive to divide the electorate. Candidates propose different policies to appeal to different voter groups, build a voter base, and secure political relevance (and rents). This strategy is more effective the more the society is fractured into two factions with very different policy stances. As a result, parties benefit from polarizing voters and increasing the ideological cohesion of opposite political groups, favoring the very same types of societal conflict that social scientists are observing in the US electorate today.
Highlighting the supply-side incentives to divide voters is just a preliminary step in the important process of understanding the determinants of polarization. It certainly calls for further research. A follow-up set of empirical questions pertains to how politicians influence voter views and how effective the efforts in this direction are. Are certain policy issues – for instance those that pin down cleavages related to identity, cultural issues, and moral values – more polarizing than others, and if yes, does political advertising on these topics fracture the electorate more than advertising on economic policy issues? Are there communication strategies that reinforce the ideological coherence of opposite political groups, resulting in a growing multidimensionality of voter conflict?
From a theoretical standpoint, it seems important to unveil the psychological mechanisms and biases in information processing that make voters vulnerable to the polarizing effects of the communication strategies of parties and partisan information sources. Reaching a deeper understanding of how the interaction between political communication and voter information processing biases might result in increased extremism, polarization, or backlash, not only seems an important avenue of research, but also a necessary step for those who wish to design interventions to curb these polarizing forces.
Appendix A Appendix
A.1 Proofs
Proof of Proposition 2.1
We prove the statement by contradiction. Assume that there exists a pure strategy equilibrium $x^{\star}\in\mathbb{R}^{2k}$ such that $x^{\star}_{A}=x^{\star}_{B}$. Since the electorate is diverse (not homogeneous), there exists a policy issue $k$ such that $x^{k}_{j}\neq x^{k}_{i}$ for some $i,j\in I$. Fix $k$. Without loss of generality, assume that voter types are indexed in such a way that $x^{k}_{1}\leq x^{k}_{2}\leq...\leq x^{K}_{N}$. Note that at least one inequality must hold strictly. Consider now deviation $x_{A}$ for party $A$, with $x^{k}_{A}>x^{k\star}_{A}$ and $x^{k^{\prime}}_{A}=x^{k^{\prime}\star}_{A}$ for each $k^{\prime}\neq k$. It is easy to see that, at action profile $(x_{A},x_{B}^{\star})$, the expected utility of party $A$ is
$V_{A}(x_{A},x_{B}^{\star})=\nu(0)+\sum_{i=1}^{N}\pi_{i}(x_{A},x_{B}^{\star})\bar{w}_{i}$
where, for each $i$, $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\bar{s}_{i}=\sum_{j\geq i}s_{i}$, with $\bar{s}_{N+1}=0$. The partial derivative of the expected utility with respect to $x_{A}^{k}$ is equal to
$\frac{\partial V_{A}(x_{A},x^{\star}_{B})}{\partial x^{k}_{A}}=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x_{A},x_{B}^{\star})}{\partial x^{k}_{A}}\bar{w}_{i}$
and letting $x^{k}_{A}$ approach $x^{k\star}_{A}$ from above, the right derivative with respect to $x^{k}_{A}$ at $x^{\star}$ is
$\lim_{x^{k}_{A}\rightarrow x^{k\star+}_{A}}\frac{\partial V_{A}(x_{A},x^{\star}_{B})}{\partial x^{k}_{A}}=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\bar{w}_{i}.$
Consider now the left derivative. Formally, let $x_{A}$ be such that $x^{k}_{A}>x^{k\star}_{A}$ and $x^{k^{\prime}}_{A}=x^{k^{\prime}\star}_{A}$ for each $k^{\prime}\neq k$. By following steps similar to the ones above, it is easy to see that
$\lim_{x^{k}_{A}\rightarrow x^{k\star-}_{A}}\frac{\partial V_{A}(x_{A},x^{\star}_{B})}{\partial x^{k}_{A}}=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\underaccent{\bar}{w}_{i}.$
where, for each $i$, $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$.
We now show that $\lim_{x^{k}_{A}\rightarrow x^{k\star-}_{A}}\frac{\partial V_{A}(x_{A},x^{\star}_{B})}{\partial x^{k}_{A}}-\lim_{x^{k}_{A}\rightarrow x^{k\star+}_{A}}\frac{\partial V_{A}(x_{A},x^{\star}_{B})}{\partial x^{k}_{A}}<0$, which proves that $x^{\star}_{A}$ is not a best response to $x^{\star}_{B}$. Note that the difference between limits is equal to
$\sum_{i=1}^{N}\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}(\underaccent{\bar}{w}_{i}-\bar{w}_{i})=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\left[(\nu(1-\bar{s}_{i+1})+\nu(\bar{s}_{i+1}))-(\nu(1-\bar{s}_{i})+\nu(\bar{s}_{i}))\right]$
$=\left(\frac{\partial\pi_{N}(x^{\star})}{\partial x^{k}_{A}}-\frac{\partial\pi_{1}(x^{\star})}{\partial x^{k}_{A}}\right)(\nu(1)+\nu(0))+\sum_{i=2}^{N}\left(\frac{\partial\pi_{i-1}(x^{\star})}{\partial x^{k}_{A}}-\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\right)\left(\nu(1-\bar{s}_{i})+\nu(\bar{s}^{i})\right).$
Next, note that, given the concavity of $v^{k}(x^{k}_{A},x_{i}^{k})$ in $|x^{k}_{A}-x_{i}|$ and $\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}=f(0)\frac{\partial v^{k}(x^{k}_{A},x_{i}^{k})}{\partial x_{A}^{k}}$, it has
$\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\geq\frac{\partial\pi_{i-1}(x^{\star})}{\partial x^{k}_{A}}\text{ for }i=2,...,N$
which holds strictly for at least some $i\in I$. Finally, using remark 1, it holds that $\nu(1-\bar{s}_{i})+\nu(\bar{s}_{i})>\nu(1)+\nu(0)$ for $i\geq 2$. It follows that
$\sum_{i=1}^{N}\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}(\omega_{i}-w_{i})<\underbrace{\left[\frac{\partial\pi_{N}(x^{\star})}{\partial x^{k}_{A}}-\frac{\partial\pi_{1}(x^{\star})}{\partial x^{k}_{A}}+\sum_{i=2}^{N}\left(\frac{\partial\pi_{i-1}(x^{\star})}{\partial x^{k}_{A}}-\frac{\partial\pi_{i}(x^{\star})}{\partial x^{k}_{A}}\right)\right]}_{\begin{subarray}{c}=0\end{subarray}}\left(\nu(1)+\nu(0)\right)$
which proves that $x_{A}^{\star}$ is not a best response to $x_{B}^{\star}$. Hence $x^{\star}$ is not a Nash equilibrium. $\blacksquare$
Proof of Proposition 3.1
By proposition 2.1, we know that for any platform $x\in\mathbb{R}$, $x$ is never the best response to $x$. Fix a distribution $(\mathbf{x},\mathbf{s})$ and order $i,...,N$ so that $x_{1}\leq x_{2}\leq...\leq x_{N}$.
Fix $p\in\{A,B\}$ and let $x_{-p}\in\mathbb{R}$. As a first step we show that, if $x_{p}$ is a best response to $x_{-p}$, then either $x_{p}>x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$, for $\bar{w}_{i}=\frac{\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})}{\nu(1)-\nu(0)}$, or $x_{p}<x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$, for $\underaccent{\bar}{w}_{i}=\frac{\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})}{\nu(1)-\nu(0)}$.
Assume that $x_{p}$ is a best response to $x_{-p}$. Given that $x_{p}\neq x_{-p}$ there are two cases. The first is $x_{p}>x_{-p}$. Subject to the constraint $x_{p}>x$, and fixing $x_{-p}$,$V_{p}$ is strictly concave in $p$’s platform $x$. Hence, $x_{p}$ it must satisfy the first-order condition $\frac{\partial V_{p}}{\partial x_{p}}=0$. When $p=A$, the first order condition is
$\frac{\partial V_{A}(x_{A},x_{B})}{\partial x_{A}}=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x_{A},x_{B})}{\partial x_{A}}\bar{w}_{i}=0.$ (5)
Note that $\pi_{i}(x_{A},x_{B})=\frac{1}{2}+\frac{\Delta_{i}}{2\phi}$ due to the uniform distribution assumption, so it has
$\frac{\partial\pi_{i}(x_{A},x_{B})}{\partial x_{A}}=\frac{x_{i}-x_{A}}{\phi}\text{ for }i=1,...,N.$
Plugging the above expression in 5, we obtain,
$\frac{\partial V_{A}(x_{A},x_{B})}{\partial x_{A}}=0\iff\sum_{i=1}^{N}\frac{x_{i}-x_{A}}{\phi}\bar{w}_{i}=0\iff x_{A}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$
Next, consider the second case, when the best response to $x_{-p}$ is strictly less than $x_{p}$. Because of strict quasi-concavity of $V_{p}$ in party $p$’s platform $x$ under the constrain $x<x_{-p}$, $x_{p}$ must satisfy the first order condition $\frac{\partial V_{p}}{\partial x_{p}}=0$. When $p=A$, this requires
$\frac{\partial V_{A}(x_{A},x_{B})}{\partial x_{A}}=\sum_{i=1}^{N}\frac{\partial\pi_{i}(x_{A},x_{B})}{\partial x_{A}}\underaccent{\bar}{w}_{i}=0.$ (6)
Substituting for $\frac{\partial\pi_{i}(x_{A},x_{B})}{\partial x_{A}}=\frac{x_{i}-x_{A}}{\phi}$ in 6, we obtain,
$\frac{\partial V_{A}(x_{A},x_{B})}{\partial x_{A}}=0\iff\sum_{i=1}^{N}\frac{x_{i}-x_{A}}{\phi}\underaccent{\bar}{w}_{i}=0\iff x_{A}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}.$
Note that we could follow the exact same steps for $p=B$, and, as parties are identical with respect of the primitives of the game, we would indeed find the same best responses, $x_{p}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$ if $x_{p}>x_{-p}$, and $x_{p}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$ if $x_{p}<x_{-p}$. Hence, the game admits at most two best responses, $\underaccent{\bar}{x}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$ and $\bar{x}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$.
In order to show that a pure strategy Nash equilibrium exists, we need to show that $\bar{x}$ and $\underaccent{\bar}{x}$ are the best responses to each other. It suffices to show that $\underaccent{\bar}{x}<\bar{x}$, as we argue below. First, by proposition 2.1, $\underaccent{\bar}{x}$ is not the best response to $\underaccent{\bar}{x}$ and $\bar{x}$ is not the best response to $\bar{x}$. Hence, by the previous step of the proof, $\underaccent{\bar}{x}$ is the only possible best response to $\bar{x}$ and $\bar{x}$ is the only possible best response to $\underaccent{\bar}{x}$. Second, note that, for each $p=A,B$, $V_{p}$ is continuous, increasing in $x_{p}$ when $x_{p}<\underaccent{\bar}{x}$, decreasing in $x_{p}$ when $x_{p}>\bar{x}$. Hence, given the strict concavity of $V_{p}$ in $x_{p}$ when $x_{p}\neq x_{-p}$, $x_{p}$ is the best response to $\underaccent{\bar}{x}$ if $x_{p}>\underaccent{\bar}{x}$ and it satisfies the FOC of the respective maximization problem. The FOC requires $x_{p}=\bar{x}$. Similarly, $x_{p}$ is the best response to $\bar{x}$ if $x_{p}<\bar{x}$ and it satisfies the FOC of the respective maximization problem. The FOC requires $x_{p}=\underaccent{\bar}{x}$. Hence, $\underaccent{\bar}{x}<\bar{x}$ is a sufficient condition for $\bar{x}$ and $\underaccent{\bar}{x}$ to be the best responses to each other.
We now show $\underaccent{\bar}{x}<\bar{x}$. First, note that since $\underaccent{\bar}{w}_{i},\bar{w}_{i}\in[0,1]$ for $i=1,...,N$, and $\sum_{i=1}^{N}\bar{w}_{i}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=1$, $(\bar{w}_{i},x_{i})_{i\in I}$ and $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$ are probability distributions.
Second, $(\bar{w}_{i},x_{i})_{i\in I}$ strictly first order dominates $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$, as we now show. Since the probability distributions share support $\{x_{1},...,x_{N}\}$, it is sufficient to show that, for each $i=1,...,N$
$\sum_{j\geq i}^{N}\bar{w}_{j}\geq\sum_{j\geq i}^{N}\underaccent{\bar}{w}_{j}$
with the inequality holding strictly for at least some $i$. Fix $i\in I$, it holds
$\sum_{j\geq i}^{N}\bar{w}_{j}=\sum_{j\geq i}^{N}(\nu(\bar{s}_{j})-\nu(\bar{s}_{j+1}))$ $=\nu(\bar{s}_{i})-\nu(0)$
$\geq\nu(1)-\nu(1-\bar{s}_{i})=\sum_{j\geq i}^{N}(\nu(1-\bar{s}_{j+1})-\nu(1-\bar{s}_{j}))=\sum_{j\geq i}^{N}\underaccent{\bar}{w}_{j}$
where the inequality hold strictly for every $i>1$.
Having proved strict first order stochastic dominance, note that $\bar{x}$ is the expected value of $(\bar{w}_{i},x_{i})_{i\in I}$ and $\underaccent{\bar}{x}$ is the expected value of $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$. Strict first order stochastic dominance implies $\bar{x}>\underaccent{\bar}{x}$. This concludes the proof that the pure strategy equilibrium exists, and that only $\bar{x}$ and $\underaccent{\bar}{x}$ will be played in such equilibrium.
Next, note that for each $i\in I$, it has $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$. The former weight represents the payoff gain of the party offering $\bar{x}$ when she wins type $i$. Voter group $i$ will vote for the party offering $\bar{x}$ only if all $j>i$ – that is, a share $\bar{s}_{i+1}$ of the distribution, also vote for that party. If voter group $i$ also votes for the party offering $\bar{x}$, the vote share of that party increases from $\bar{s}_{i+1}$ to $\bar{s}_{i}=\bar{s}_{i+1}+s_{i}$, with an increase in payoff equal to $\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})=\bar{w}_{i}$. Using the same argument, it is easy to show that $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$ is the payoff gain of the party offering $\underaccent{\bar}{x}$ from winning voter group $i$. Note that, by remark 1, it has
$\bar{s}_{i}<\frac{1}{2}\implies\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})>\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$
$\bar{s}_{i}>\frac{1}{2}\implies\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})<\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$
which proves that $\bar{w}_{i}>\underaccent{\bar}{w}_{i}$ if $x_{i}>x_{m}$ and $\bar{w}_{i}<\underaccent{\bar}{w}_{i}$ if $x_{i}<x_{m}$. Finally, assume, without loss of generality that $x^{A}=\bar{x}$ in equilibrium. It has
$V_{A}(\bar{x},\underaccent{\bar}{x})=\nu(0)+\sum_{i=1}^{N}\pi_{i}(\bar{x},\underaccent{\bar}{x})\bar{w}_{i}$ $=\frac{\nu(1)+\nu(0)}{2}+\sum_{i=1}^{N}\frac{\underaccent{\bar}{x}^{2}-\bar{x}^{2}-2x_{i}(\underaccent{\bar}{x}-\bar{x})}{2\phi}\bar{w}_{i}$
$=\frac{\nu(1)+\nu(0)}{2}+\frac{(\bar{x}-\underaccent{\bar}{x})^{2}}{2\phi}$
and
$V_{B}(\bar{x},\underaccent{\bar}{x})=\nu(1)-\sum_{i=1}^{N}\pi_{i}(\bar{x},\underaccent{\bar}{x})\omega_{i}$ $=\frac{\nu(1)+\nu(0)}{2}+\sum_{i=1}^{N}\frac{\bar{x}^{2}-\underaccent{\bar}{x}^{2}-2x_{i}(\bar{x}-\underaccent{\bar}{x})}{2\phi}\underaccent{\bar}{w}_{i}$
$=\frac{\nu(1)+\nu(0)}{2}+\frac{(\bar{x}-\underaccent{\bar}{x})^{2}}{2\phi},$
which concludes the proof of the proposition $\blacksquare$
Proof of Proposition 3.2
Fix $(\mathbf{x},\mathbf{s})=(x_{i},s_{i})_{i\in I}$, and order indexes such that $x_{1}\leq x_{2}\leq...\leq x_{N}$. Recall that $\hat{V}_{p}(\mathbf{x},\mathbf{s})=\hat{V}_{p}(\bar{x},\underaccent{\bar}{x})$ by definition, hence $\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}\propto\bar{w}_{i}-\underaccent{\bar}{w}_{i}$, which is strictly positive if $x_{i}>x_{m}$ and strictly negative if $x_{i}<x_{m}$, by proposition 3.1. It follows that
$\frac{\partial\pi_{i}(\mathbf{x},\mathbf{s})}{\partial x_{i}}=\frac{(\underaccent{\bar}{x}\underaccent{\bar}{w}_{i}-\bar{x}\bar{w}_{i})+(\bar{x}-\underaccent{\bar}{x})+x_{i}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})}{\phi}$
Note that if $x_{i}>\max\{\underaccent{\bar}{x},x_{m}\}$ then
$(\underaccent{\bar}{x}\underaccent{\bar}{w}_{i}-\bar{x}\bar{w}_{i})+(\bar{x}-\underaccent{\bar}{x})+x_{i}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})$ $>(\underaccent{\bar}{x}\underaccent{\bar}{w}_{i}-\bar{x}\bar{w}_{i})+(\bar{x}-\underaccent{\bar}{x})+\underaccent{\bar}{x}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})$
$=(1-\bar{w}_{i})(\bar{x}-\underaccent{\bar}{x})>0,$
and, given that $x_{i}>x_{m}\implies\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}>0$, it has $\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}\frac{\partial\pi_{i}(\mathbf{x},\mathbf{s})}{\partial x_{i}}>0$.
Similarly, if $x_{i}<\min\{\bar{x},x_{m}\}$,
$(\underaccent{\bar}{x}\underaccent{\bar}{w}_{i}-\bar{x}\bar{w}_{i})+(\bar{x}-\underaccent{\bar}{x})$ $+x_{i}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})=(1-\bar{w}_{i})\bar{x}-(1-\underaccent{\bar}{w}_{i})\underaccent{\bar}{x}+x_{i}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})$
$>(1-\bar{w}_{i})\bar{x}-(1-\underaccent{\bar}{w}_{i})\underaccent{\bar}{x}-(\underaccent{\bar}{w}_{i}-\bar{w}_{i})\underaccent{\bar}{x}=(1-\bar{w}_{i})(\bar{x}-\underaccent{\bar}{x})>0$
and, given that $x_{i}<x_{m}\implies\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}<0$, it has $\frac{\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}\frac{\partial\pi_{i}(\mathbf{x},\mathbf{s})}{\partial x_{i}}<0$. By the same argument, it is easy to show that if $x_{i}<\min\{\bar{x},{x_{m}}\}$ then $\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}\frac{\partial(1-\pi_{i}(\mathbf{x},\mathbf{s}))}{\partial x_{i}}>0$ and if $x_{i}>\max\{\bar{x},{x_{m}}\}$ then $\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}\frac{\partial(1-\pi_{i}(\mathbf{x},\mathbf{s}))}{\partial x_{i}}<0$, concluding the proof. Since $\underaccent{\bar}{w}_{i},\bar{w}_{i}\in(0,1)$ for each $i=1,...,N$, we have $x_{1}<\underaccent{\bar}{x}$ and $x_{N}>\bar{x}$, whenever $x_{m}\notin\{x_{1},x_{N}\}$ which proves the remark $\blacksquare$
Proof of Proposition 3.3
[…]
Next, note that
$\bar{x}-\underaccent{\bar}{x}=\sum_{i\in I}(\bar{w}_{i}-\underaccent{\bar}{w}_{i})x_{i}=\sum_{i\in I}[(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))]x_{i}$
$\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}=\sum_{i\in I^{\prime}}(\bar{w}_{i}^{\prime}-\underaccent{\bar}{w}_{i}^{\prime})x_{i}^{\prime}=\sum_{i\in I^{\prime}}[(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))]x^{\prime}_{i}.$
Consider distributions $(\mathbf{x}^{\prime}_{R},\mathbf{s}^{\prime}_{R})$ and $(\mathbf{x}_{R},\mathbf{s}_{R})$, and recall that $(\mathbf{x}^{\prime}_{R},\mathbf{s}^{\prime}_{R})$ first-order stochastically dominates $(\mathbf{x}^{\prime}_{R},\mathbf{s}^{\prime}_{R})$. Next, we show that
$\left(x^{\prime}_{Ri},\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\right)_{i\in I^{\prime}_{R}}$
first order stochastically dominates
$\left(x_{Ri},\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\right)_{i\in I_{R}}.$
For simplicity, in the proof we assume that the median type $x_{m}=0$ is such that exactly $50\%$ of the two voter distributions considered fall to the left and to the right of the median type (implying that no voter has the median type). [^27] The proof for the general case follows the same steps as the one presented below, except that when the median type has non-zero mass, then it must be split in two “virtual” types, $x_{m+}$, $x_{m-}$, with $x_{m+}=x_{m-}=x_{m}=0$, $s_{m-}=\frac{1}{2}-\sum_{i:x_{i}<0}s_{i}$, $s_{m+}=\frac{1}{2}-\sum_{i:x_{i}>0}s_{i}$. When defining the conditional distribution $(\mathbf{x}_{R},\mathbf{s}_{R})$, type $x_{m+}$ is included in the support as the lowest type, with probability $2s_{m+}$. When defining the conditional distribution$(\mathbf{x}_{L},\mathbf{s}_{L})$, type $x_{m-}$ is included in the support as the highest type, with probability $2s_{m-}$. Type $x_{m}$ is not included in any of the two distributions
Consider any $\hat{x}\in supp(\mathbf{x}^{\prime}_{R},\mathbf{s}^{\prime}_{R})\cup supp(\mathbf{x}_{R},\mathbf{s}_{R})$. If $\hat{x}>x_{|I|}$, define $x_{|I|+1}=\hat{x}$ and $\bar{s}_{|I|+m}=0$ for $m=1,2$. Fix $i\in I$ and $j\in I^{\prime}$ such that $x_{i-1}<\hat{x}\leq x_{i}$ and $x^{\prime}_{j-1}<\hat{x}\leq x^{\prime}_{j}$.
Note that by first order stochastic dominance, $\bar{s}^{\prime}_{Rj}\geq\bar{s}_{Ri}$, which implies $\bar{s}^{\prime}_{j}\geq\bar{s}_{i}$. Next, note that
$\sum_{n:x_{n}\geq x_{i}}\frac{(\nu(\bar{s}_{n})+\nu(1-\bar{s}_{n}))-(\nu(\bar{s}_{n+1})+\nu(1-\bar{s}_{n+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}=\frac{\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i})-(\nu(1)+\nu(0))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$
$\sum_{n:x^{\prime}_{n}\geq x^{\prime}_{j}}\frac{(\nu(\bar{s}^{\prime}_{n})+\nu(1-\bar{s}^{\prime}_{n}))-(\nu(\bar{s}^{\prime}_{n+1})+\nu(1-\bar{s}^{\prime}_{n+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}=\frac{\nu(\bar{s}^{\prime}_{j})+\nu(1-\bar{s}^{\prime}_{j})-(\nu(1)+\nu(0))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$
and that $\frac{\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i})-(\nu(1)+\nu(0))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\leq\frac{\nu(\bar{s}^{\prime}_{j})+\nu(1-\bar{s}^{\prime}_{j})-(\nu(1)+\nu(0))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$ since $\frac{1}{2}\geq\bar{s}^{\prime}_{j}\geq\bar{s}_{i}$, which proves our statement that $\left(x^{\prime}_{Ri},\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\right)_{i\in I^{\prime}_{R}}$ first order stochastically dominates $\left(x_{Ri},\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\right)_{i\in I_{R}}$.
Consider now distributions $(\mathbf{x}^{\prime}_{L},\mathbf{s}^{\prime}_{L})$ and $(\mathbf{x}_{L},\mathbf{s}_{L})$, and recall that $(\mathbf{x}^{\prime}_{L},\mathbf{s}^{\prime}_{L})$ is first-order stochastically dominated by $(\mathbf{x}_{L},\mathbf{s}_{L})$. We show that
$\left(x^{\prime}_{Li},\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}\right)_{i\in I^{\prime}_{L}}$
is first order stochastically dominated by
$\left(x_{Li},\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}\right)_{i\in I_{L}}.$
Consider any $\hat{x}\in supp(\mathbf{x}^{\prime}_{L},\mathbf{s}^{\prime}_{L})\cup supp(\mathbf{x}_{L},\mathbf{s}_{L})$. If $\hat{x}<x^{\prime}_{1}$, let $x_{0}<\hat{x}$ and $s_{0}=0$. Fix $i\in I$ and $j\in I^{\prime}$ such that $x_{i-1}<\hat{x}\leq x_{i}$ and $x^{\prime}_{j-1}<\hat{x}\leq x^{\prime}_{j}$. Note that by first order stochastic dominance, $\bar{s}^{\prime}_{Lj}\leq\bar{s}_{Li}$, which implies $\bar{s}^{\prime}_{j}\leq\bar{s}_{i}$. Next, note that
$\sum_{n\geq i}^{|I_{L}|}\frac{(\nu(\bar{s}_{n})+\nu(1-\bar{s}_{n}))-(\nu(\bar{s}_{n+1})+\nu(1-\bar{s}_{n+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}=\frac{2\nu(\frac{1}{2})-(\nu(\bar{s}_{n})+\nu(1-\bar{s}_{n}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$
$\sum_{n\geq j}^{|I^{\prime}_{L}|}\frac{(\nu(\bar{s}^{\prime}_{n})+\nu(1-\bar{s}^{\prime}_{n}))-(\nu(\bar{s}^{\prime}_{n+1})+\nu(1-\bar{s}^{\prime}_{n+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}=\frac{2\nu(\frac{1}{2})-(\nu(\bar{s}^{\prime}_{n})+\nu(1-\bar{s}^{\prime}_{n}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$
and that $\frac{2\nu(\frac{1}{2})-(\nu(\bar{s}_{n})+\nu(1-\bar{s}_{n}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}\geq\frac{2\nu(\frac{1}{2})-(\nu(\bar{s}^{\prime}_{n})+\nu(1-\bar{s}^{\prime}_{n}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}$ since $\frac{1}{2}\leq\bar{s}^{\prime}_{i}\leq\bar{s}_{i}$, which proves our statement that $\left(x^{\prime}_{Li},\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}\right)_{i\in I^{\prime}_{L}}$ is first order stochastically dominated by $\left(x_{Li},\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}\right)_{i\in I_{L}}$.
These first order stochastic dominance results imply
$\sum_{i\in I_{R}}\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}x_{i}$
$\leq\sum_{i\in I^{\prime}_{R}}\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{2\nu(\frac{1}{2})-(\nu(1)+\nu(0))}x^{\prime}_{i}$
$\implies\sum_{i\in I_{R}}(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))x_{i}$
$\leq\sum_{i\in I^{\prime}_{R}}(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))x^{\prime}_{i}$
and
$\sum_{i\in I_{L}}\frac{(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}x_{i}$
$\geq\sum_{i\in I^{\prime}_{L}}\frac{(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))}{(\nu(1)+\nu(0))-2\nu(\frac{1}{2})}x^{\prime}_{i}$
$\implies\sum_{i\in I_{L}}(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))x_{i}$
$\leq\sum_{i\in I^{\prime}_{L}}(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))x^{\prime}_{i}$
[…]
$\sum_{i\in I}\left[(\nu(\bar{s}_{i})+\nu(1-\bar{s}_{i}))-(\nu(\bar{s}_{i+1})+\nu(1-\bar{s}_{i+1}))\right]x_{i}$
$<$ $\sum_{i\in I^{\prime}}\left[(\nu(\bar{s}^{\prime}_{i})+\nu(1-\bar{s}^{\prime}_{i}))-(\nu(\bar{s}^{\prime}_{i+1})+\nu(1-\bar{s}^{\prime}_{i+1}))\right]x^{\prime}_{i},$
which is equivalent to $\bar{x}-\underaccent{\bar}{x}<\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}$. Finally, recall that
$V_{p}(\bar{x},\underaccent{\bar}{x})=\frac{\nu(1)+\nu(0)}{2}+\frac{(\bar{x}-\underaccent{\bar}{x})^{2}}{2\phi}$
for $p=A,B$. Since $0<\bar{x}-\underaccent{\bar}{x}<\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}$, both parties are strictly better off, in equilibrium, under $(\textbf{x}^{\prime},\textbf{s}^{\prime})$ than under $(\textbf{x},\textbf{s})$blacksquare
Proof of Lemma 4.1
Before starting the proof of the lemma, it is useful to make two preliminary observations.
First, we prove that if $x\in\mathbb{R}^{2K}$ is a local equilibrium profile, then there cannot be $i,j\in I$, $i\neq j$, such that $\Delta_{i}(x)=\Delta_{j}(x)$. The proof is by contradiction. Assume that there is $x=(x_{A},x_{B})\in\mathcal{L}$ such that $\exists i,j\in I,i\neq j$ with $\Delta_{i}(x)=\Delta_{j}(x)$. Since $x_{i}\neq x_{j}$, one can find $k\leq K$ such that $x^{k}_{i}\neq x^{k}_{j}$. Fix such $k$. Now consider any subset $G\subset I$ such that (i) for each $i,j\in G$, $\Delta_{i}(x)=\Delta_{j}(x)$ and (ii) $\exists i,j\in G$ such that $x_{i}^{k}\neq x_{j}^{k}$. Note that $G$ exists by construction. Next, let $R\in\mathcal{R}(G)$ be any order over types in $G$ such that for every two positions $i,j\in\{1,2,...,|G|\}$, $i>j\implies x^{k}_{R_{i}}\geq x^{k}_{R_{j}}$, where $R_{j}$ denotes the type in position $j$ of the ranking. Next, let $\Delta=\max_{i\in G}\Delta_{i}(x)$, and define $\bar{s}=\sum_{j\in I:\Delta_{j}(x)>\Delta}s_{j}$ to be the total share of voters that are more likely to vote $A$ than voters in $G$. Define, for each $i=1,...,|G|$, $\bar{s}_{R_{i}}=\bar{s}+\sum_{j=i}^{|G|}s_{R_{j}}$. We will show that
$\lim_{\hat{x}^{k}_{A}\rightarrow x^{k+}_{A}}\sum_{i=1}^{|G|}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{+}-\lim_{\hat{x}^{k}_{A}\rightarrow x^{k-}_{A}}\sum_{i=1}^{|G|}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{-}>0,$
where $\bar{w}_{R_{i}}^{+}=\nu(\bar{s}_{R_{i}})-\nu(\bar{s}_{R_{i+1}})$ and $\bar{w}_{i}^{-}=\nu(\bar{s}+\bar{s}_{R_{1}}-\bar{s}_{R_{i+1}})-\nu(\bar{s}+\bar{s}_{R_{1}}-\bar{s}_{R_{i}})$, and $\bar{s}_{R_{|G|+1}}=\bar{s}$. Note that the above difference of limits is equal to
$\sum_{i=1}^{|G|}\bar{w}^{+}_{R_{i}}x^{k}_{R_{i}}-\sum_{i=1}^{|G|}\bar{w}^{-}_{R_{i}}x^{k}_{R_{i}}>0$
which is strictly greater than 00 because $(x^{k}_{Ri},\frac{\bar{w}_{R_{i}}^{+}}{\nu(\bar{s}_{1})-\nu(\bar{s})})_{i=1}^{|G|}$ strictly first order stochastically dominates $(x^{k}_{R_{i}},\frac{\bar{w}_{R_{i}}^{-}}{\nu(\bar{s}_{1})-\nu(\bar{s})})_{i=1}^{|G|}$ by the strict concavity of $\nu$.
Next, we show that for any $i\in I$ such that $\nexists j,\Delta_{i}(x)=\Delta_{j}(x),x^{k}_{i}\neq x^{k}_{j}$, then the relative position of type $i$ in the ordering of types in $I$ based on $\Delta_{i}(x)$ does not change for small changes in $x_{A}^{k}$. Let $R\in\mathcal{R}$ be a (strict) type ordering such that $i>j\implies\Delta_{R_{i}}(x)\geq\Delta_{R_{j}}(x)$. Let $i\in\{1,...,N\}$ be such that $\nexists j\in\{1,...,N\},j\neq i$ such that $\Delta_{R_{i}}(x)=\Delta_{R_{j}}(x)$ and $x^{k}_{R_{i}}\neq x^{k}_{R_{j}}$. It has
$\lim_{\hat{x}^{k}_{A}\rightarrow x^{k+}_{A}}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{+}-\lim_{\hat{x}^{k}_{A}\rightarrow x^{k-}_{A}}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{-}=0,$
where $\bar{w}^{+}_{R_{i}}=\bar{w}^{-}_{R_{i}}=\nu(\bar{s}_{R_{i}})-\nu(\bar{s}_{R_{i+1}})=\bar{w}_{R_{i}}$. To see why $w_{R_{i}}$ is the same for the right and left limits, note that it only depends on $\bar{s}_{R_{i}}=\bar{s}_{R_{i+1}}+s_{R_{i}}$ and $\bar{s}_{R_{i+1}}=\sum_{j=i+1}^{|I|}s_{R_{j}}$. By construction $\nexists j\in\{1,...,|I|\},j\neq i$ such that $\Delta_{R_{i}}(x)=\Delta_{R_{j}}(x)$ and $x^{k}_{R_{i}}\neq x^{k}_{R_{j}}$, so $\bar{s}_{R_{i+1}}$ is constant in a neighborhood of $x_{A}^{k}$: the set of types in positions strictly above $R_{i}$ is unaffected by small enough changes in $x_{A}^{k}$.
It follows that
$\lim_{\hat{x}^{k}_{A}\rightarrow x^{k+}_{A}}\frac{\partial V_{A}(x)}{\partial x^{k}_{A}}-\lim_{\hat{x}^{k}_{A}\rightarrow x^{k-}_{A}}\frac{\partial V_{A}(x)}{\partial x^{k}_{A}}$
$=$ $\lim_{\hat{x}^{k}_{A}\rightarrow x^{k\star+}_{A}}\sum_{i=1}^{|I|}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{+}-\lim_{\hat{x}^{k}_{A}\rightarrow x^{k-}_{A}}\sum_{i=1}^{|I|}\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}^{-}>0$
which proves that party $A$ has a profitable marginal deviation from $x^{k}_{A}$. Hence $x\notin\mathcal{L}$. This proves that if $x\in\mathcal{L}$, then for each $i,j\in\mathcal{L},i\neq j$, it has $\Delta_{i}(x)\neq\Delta_{j}(x)$.
Now, let $x\in\mathcal{L}$, and let $R\in R(I)$ represent the strict ordering of types in $I$ such that $i>j\implies\Delta_{i}(x)>\Delta_{j}(x)$. First, such an ordering is well-defined because of our previous result. Second, since the ordering is strict, it is preserved in a neighborhood of $x$: because of the continuity of $\Delta_{i}(x)$, $\Delta_{R_{1}}(x)<\Delta_{R_{2}}(x)<...<\Delta_{R_{|I|}}(x)$ implies $\Delta_{i}(x^{\prime})$, $\Delta_{R_{1}}(x^{\prime})<\Delta_{R_{2}}(x^{\prime})<...<\Delta_{R_{|I|}}(x^{\prime})$ for $x^{\prime}$ in a neighborhood of $x$. For each $i\in\{1,...,|I|\}$, let $\bar{s}_{R_{i}}=\sum_{j=i}^{|I|}s_{R_{i}}$ and $\bar{w}_{R_{i}}=\nu(\bar{s}_{R_{i}})-\nu(\bar{s}_{R_{i+1}})$. The above implies that, for each $i,k,k^{\prime},p,p^{\prime}$, $\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}\bar{w}_{R_{i}}$ and $\frac{\partial^{2}\pi_{R_{i}}(x)}{\partial x^{k}_{p}\partial x^{k^{\prime}}_{p^{\prime}}}\bar{w}_{R_{i}}$ exist and have the same sign as $\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}$ and $\frac{\partial^{2}\pi_{R_{i}}(x)}{\partial x^{k}_{A}\partial x^{k^{\prime}}_{p^{\prime}}}$. In particular, given our assumptions on payoffs, $\frac{\partial^{2}\pi_{R_{i}}(x)}{\partial x^{k}_{A}\partial x^{k^{\prime}}_{p^{\prime}}}<0$ if $A=p^{\prime}$ and $k=k^{\prime}$, and $\frac{\partial^{2}\pi_{R_{i}}(x)}{\partial x^{k}_{A}\partial x^{k^{\prime}}_{p^{\prime}}}=0$ otherwise. Because $x\in\mathcal{L}$ and the SOC binds in a neighborhood of $x$, the FOC must also bind. It follows that $\frac{\partial\pi_{R_{i}}(x)}{\partial x^{k}_{A}}=0$, and, $\frac{\partial V_{A}(x)}{\partial x^{k}_{A}}=0$ for each $k$, $\frac{\partial^{2}V_{A}(x)}{\partial x^{k}_{A}\partial x^{k}_{A}}<0$, and $\frac{\partial^{2}V_{A}(x)}{\partial x^{k}_{A}\partial x^{k^{\prime}}_{p^{\prime}}}=0$ for $k^{\prime}\neq k$ or $p^{\prime}\neq A$. In a similar fashion, on can show that $\frac{\partial^{2}V_{B}(x)}{\partial x^{k}_{B}}=0$ for each $k$, $\frac{\partial^{2}V_{B}(x)}{\partial x^{k}_{B}\partial x^{k}_{A}}<0$, and $\frac{\partial^{2}V_{B}(x)}{\partial x^{k}_{B}\partial x^{k^{\prime}}_{p^{\prime}}}=0$ for $k^{\prime}\neq k$ or $p^{\prime}\neq B$
We have shown that every local equilibrium $x=(x_{A},x_{B})$ must induce a strict ordering over types, based on their $(\Delta_{i}(x))_{i\in I}$, and that the payoff function $V_{p}$ is differentiable at any profile inducing such a strict ordering. Hence, for each $p=A,B$, $x_{p}$ must satisfy the first order condition $\frac{\partial V_{p}(x)}{\partial x^{k}_{p}}=0$, for $k=1,...,K$. But given that the payoff function is strictly concave in a neighborhood of every differentiability point, the first order condition is also sufficient for $x_{p}$ to be a local best response to $x_{-p}$.
We summarize the results in the following lemma.
Lemma A.1
A policy platform $x=(x^{k})_{k\leq K}\in\mathbb{R}^{K}$ for party $A$ is a local best response to platform $\hat{x}$ for party $B$ if and only if $\frac{\partial V_{A}(x,\hat{x})}{\partial x^{k}}=0$ for each $k=1,...,K$. A policy platform $x=(x^{k})_{k\leq K}\in\mathbb{R}^{K}$ for party $B$ is a local best response to platform $\hat{x}$ for party $A$ if and only if $\frac{\partial V_{B}(\hat{x},x)}{\partial x^{k}}=0$ for each $k=1,...,K$. Moreover, if $\hat{x}$ is a local best response to $x$ or $x$ is a local best response to $\hat{x}$ then $\Delta_{i}(x,\hat{x})\neq\Delta_{j}(x,\hat{x})$ for each $i,j\in I,i\neq j$.
We now prove lemma 4.1. Let $\hat{x}\in\mathcal{L}$. Let $R\in\mathcal{R}(I)$ be such that for each $i,j\in\{1,2,..,|I|\}$, then $i>j$ implies $\Delta_{R_{i}}(\hat{x})>\Delta_{R_{j}}(\hat{x})$. By lemma A.1, $R$ exists. Moreover, by A.1, the first order condition $\frac{\partial V_{p}(\hat{x})}{\partial x^{k}_{p}}=0$ for each $p=A,B$ and $k=1,...,K$ must hold. Note that
$\frac{\partial V_{A}(\hat{x})}{\partial x^{k}_{A}}=0$
$\iff$ $\sum_{i=1}^{|I|}\frac{\partial\pi_{R_{i}}(\hat{x})}{\partial\hat{x}^{k}_{A}}\bar{w}_{R_{i}}=0$
$\iff$ $x^{k}_{A}=\sum_{i=1}^{|I|}\bar{w}_{R_{i}}x^{k}_{R_{i}}=\bar{x}(R)$
and
$\frac{\partial V_{B}(\hat{x})}{\partial x^{k}_{B}}=0$
$\iff$ $\sum_{i=1}^{|I|}\frac{\partial\pi^{B}_{R_{i}}(\hat{x})}{\partial x^{k}_{B}}\underaccent{\bar}{w}_{R_{i}}=0$
$\iff$ $\hat{x}^{k}_{B}=\sum_{i=1}^{|I|}\underaccent{\bar}{w}_{R_{i}}x^{k}_{R_{i}}=\underaccent{\bar}{x}(R)$
where $\bar{w}_{R_{i}}=\nu(\bar{s}_{R_{i}})-\nu(\bar{s}_{R_{i+1}})$, $\underaccent{\bar}{w}_{R_{i}}=\nu(1-\bar{s}_{R_{i+1}})-\nu(1-\bar{s}_{R_{i}})$, and $\bar{s}_{R_{i}}=\sum_{j=i}^{N}s_{R_{i}}$, with $\bar{s}_{R_{|I|+1}}=0$. Hence $R\in\mathcal{r}(\bar{x}(R),\underaccent{\bar}{x}(R))$ and $\hat{x}\in\mathcal{L}_{\mathcal{R}}$. This proves $\mathcal{L}\subseteq\mathcal{L}_{\mathcal{R}}$. Note that the proof also implies $\mathcal{L}^{p}\subseteq\mathcal{L}^{p}_{\mathcal{R}}$.
Now, let $\hat{x}\in\mathcal{L}_{\mathcal{R}}$. By construction, there exists $R\in\mathcal{R}(I)$ such that $\hat{x}=(\bar{x}(R),\underaccent{\bar}{x}(R))$ and $R\in\mathcal{r}(\hat{x})$. Since $R\in\mathcal{r}(\hat{x})$ exists, $\Delta_{R_{1}}(\hat{x})<\Delta_{R_{2}}(\hat{x})<...<\Delta_{R_{|I|}}(\hat{x})$, then $V_{A}$ and $V_{B}$ are differentiable at $\hat{x}$ by the same argument of the previous lemma. Note that, for each $k=1,...,K$
$\frac{\partial V_{A}(\hat{x})}{\partial x^{k}_{A}}\propto\hat{x}_{A}-\sum_{i=1}^{|I|}\bar{w}_{R_{i}}x^{k}_{R_{i}}=\bar{x}(R)-\sum_{i=1}^{|I|}\bar{w}_{R_{i}}x^{k}_{R_{i}}=0$
$\frac{\partial V_{B}(\hat{x})}{\partial x^{k}_{B}}\propto\hat{x}_{B}-\sum_{i=1}^{|I|}\underaccent{\bar}{w}_{R_{i}}x^{k}_{R_{i}}=\underaccent{\bar}{x}(R)-\sum_{i=1}^{|I|}\underaccent{\bar}{w}_{R_{i}}x^{k}_{R_{i}}=0$
so that, by lemma A.1, platforms $\hat{x}_{A}$ and $\hat{x}_{B}$ are local best responses to each other. Hence $\hat{x}\in\mathcal{L}$. This proves that $\mathcal{L}=\mathcal{L}_{\mathcal{R}}$.
Finally, let $x\in\mathcal{L}$, and let $R\in\mathcal{r}(x)$ such that $x=(\bar{x}(R),\underaccent{\bar}{x}(R))$ simple algebra yields
$V_{p}(x)=\frac{\nu(1)+\nu(0)}{2}+\frac{||x_{A}-x_{B}||^{2}}{2\phi}=\frac{\nu(1)+\nu(0)}{2}+\frac{||\bar{x}(R)-\underaccent{\bar}{x}(R)||^{2}}{2\phi}$
which concludes the proof $\blacksquare$.
Proof of Lemma 4.2
For any $x\in\mathbb{R}^{K}$, let $\mathcal{B}(x)$ be the set of (global) best responses to $x$. The payoff function is continuous and it can be easily shown that best responses must lie in ×$\operatorname*{\times}_{k\leq K}[x^{k}_{\lor},x^{k}_{\land}]$, defined in section 4. Hence, for any $x\in\mathbb{R}^{K}$, $\mathcal{B}(x)$ is non-empty.
Next, note that for any $x,x^{\prime}\in\mathbb{R}^{K}$ whenever party $p=A,B$ plays $x$, party $-p$ plays $x^{\prime}$, and $x\in\mathcal{B}(x^{\prime})$, party $p$’s payoff is $\frac{\nu(1)+\nu(0)}{2}+\frac{||x-x^{\prime}||^{2}}{2\phi}$.
Define a best response sequence to be $\{x_{n}\}_{n=0}^{\infty}\in(\mathbb{R}^{2k})^{\mathbb{N}_{0}}$, $\{x_{n}\}_{n=0}^{\infty}=\{(x_{An},x_{Bn})\}_{n=0}^{\infty}$ with the following properties:
  1. (i)
    $(x_{A0},x_{B0})\in\mathbb{R}^{2K}$
  2. (ii)
    for $n>0$, $n$ odd: set $x_{An}\in\mathcal{B}(x_{Bn-1})$, with $x_{An}=x_{An-1}$ if $x_{An-1}\in\mathcal{B}(x_{Bn-1})$; $x_{Bn}=x_{Bn-1}$.
  3. (iii)
    for $n>0$, $n$ even: set $x_{Bn}\in\mathcal{B}(x_{An-1})$, with $x_{Bn}=x_{Bn-1}$ if $x_{Bn-1}\in\mathcal{B}(x_{An-1})$; $x_{An}=x_{An-1}$.
Note that we are requiring player $A$ to best respond first, but nothing would change was $B$ to move first.
Let $(\mathbf{s},\mathbf{x})\in\mathcal{V}$. Without loss of generality, assume that the median voter type vector is $\mathbf{x}_{m}=0$. Let $(x_{A0},x_{B0})\in\mathcal{L}$, and let $\{x_{n}\}_{n=0}^{\infty}\in(\mathbb{R}^{2k})^{\mathbb{N}_{0}}$ be a best response sequence initiated at $x_{0}=(x_{A0},x_{B0})$. If $x_{0}\in\mathcal{E}$, $||x_{An}-x_{Bn}||=||x_{A0}-x_{B0}||$ holds by construction for each $n\geq 0$. It remains to prove that, if $x_{0}\notin\mathcal{E}$, then $||x_{An}-x_{Bn}||>||x_{A0}-x_{B0}||$ for each $n>0$. First, we prove the following lemma. Let $x^{T}$ denote the transposition of vector $x$.
Lemma A.2
Let $(\mathbf{s},\mathbf{x})\in\mathcal{V}$, $(x_{A0},x_{B0})\in\mathcal{L}$, and $\{x_{n}\}_{n=0}^{\infty}\in(\mathbb{R}^{2k})^{\mathbb{N}_{0}}$ be a best response sequence initiated at $x_{0}=(x_{A0},x_{B0})$. Then
$||x_{An}||^{2}-||x_{Bn}||^{2}\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{An}^{T}x_{Bn}\quad\quad\forall n=1,3,5,...$ (7)
and
$||x_{Bn}||^{2}-||x_{An}||^{2}\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{Bn}^{T}x_{An}\quad\quad\forall n=2,4,6,...$ (8)
Moreover, if $x_{0}\notin\mathcal{E}$, then both inequalities hold strictly for $n>1$
The proof of lemma A.2 is by induction. First, we show that if 7 holds for some $n\geq 1$ odd, then 8 holds for $n+1$ and 7 holds for $n+2$. Then, we show that 7 holds for $n=1$.
Let $x_{0}\in\mathcal{L}$ and assume that 7 holds at $n\geq 1$, $n$ odd. By construction $x_{Bn+1}\in\mathcal{B}(x_{An})$, hence it must hold that $V_{B}(x_{An},x_{Bn+1})\geq V_{B}(x_{An},x_{Bn})$. Using the definitions of $V_{B}(x_{An},x_{Bn})$ and $V_{B}(x_{An},x_{Bn+1})$, and given that $x_{An}=x_{An+1}$ this requires
$||x_{An+1}-x_{Bn+1}||^{2}\geq||x_{An}||^{2}-||x_{Bn}||^{2}-2\underaccent{\bar}{x}(R)^{T}(x_{An}-x_{Bn})$
where $R\in\mathcal{R}(I)$ is the unique ordering of voters induced when $A$ is playing the best response $x_{An}\in\mathcal{B}(x_{Bn})$, that is $\mathcal{r}(x_{An},x_{Bn})=\{R\}$. By the symmetry of the voter distribution, (i) $R$ is symmetric, in the sense that $R_{j}=i\iff R_{|I|-j}=\bar{i}$ for each $j=1,...,|I|$, where recall that $\bar{i}$ is the type such that $x_{\bar{i}}=-x_{i}$; (ii) $\underaccent{\bar}{w}_{R_{j}}=\bar{w}_{R_{|I|-j}}$ for each $j=1,...,|I|$. As a consequence, $\underaccent{\bar}{x}(R)=-\bar{x}(R)=-x_{An}$. Hence
$||x_{An+1}-x_{Bn+1}||^{2}$ $\geq||x_{An}||^{2}-||x_{Bn}||^{2}+2x_{An}^{T}(x_{An}-x_{Bn})$
$\iff||x_{Bn+1}||^{2}-||x_{An+1}||^{2}$ $\geq||x_{An}||^{2}-||x_{Bn}||^{2}-2x_{An}^{T}x_{Bn}+2x_{An+1}^{T}x_{Bn+1}$
$\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{An+1}^{T}x_{Bn+1}$
where the last inequality follows from the inductive assumption that 7 holds at $n$. Next, note that $n+1$ is even, hence $n+2$ is odd, hence $V_{A}(x_{An+2},x_{Bn+2})\geq V_{B}(x_{An+1},x_{Bn+1})$. This implies
$||x_{An+2}-x_{Bn+2}||^{2}$ $\geq||x_{Bn+1}||^{2}-||x_{An+1}||^{2}-2\bar{x}(R)^{T}(x_{Bn+1}-x_{An+1})$
$\iff||x_{An+2}-x_{Bn+2}||^{2}$ $\geq||x_{Bn+1}||^{2}-||x_{An+1}||^{2}+2x_{Bn+1}^{T}(x_{Bn+1}-x_{An+1})$
$\iff||x_{An+2}||^{2}-||x_{Bn+2}||^{2}$ $\geq||x_{Bn+1}||^{2}-||x_{An+1}||^{2}-2x_{Bn+1}^{T}x_{An+1}+2x_{Bn+2}^{T}x_{An+2}$
$\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{Bn+2}^{T}x_{An+2}$
where the last inequality follows from the inductive assumption that 8 holds at $n+1$. We now show that $||x_{A1}||^{2}-||x_{B1}||^{2}\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{A1}^{T}x_{B1}$. Note that since $x_{A1}\in\mathcal{B}(x_{B0})$, we have $||x_{A1}-x_{B1}||^{2}\geq||x_{A0}-x_{B0}||^{2}$. Since $x_{0}\in\mathcal{L}$, $x_{A0}=\bar{x}(R)$ and $x_{B0}=\underaccent{\bar}{x}(R)$ for the same ordering $R\in\mathcal{R}(I)$. Hence, by the symmetry argument used earlier in the proof, $x_{A0}=\bar{x}(R)=-\underaccent{\bar}{x}(R)=-x_{B0}$. This implies
$||x_{A1}-x_{B1}||^{2}$ $\geq||x_{A0}-x_{B0}||^{2}$
$\iff||x_{A1}-x_{B1}||^{2}$ $\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+\frac{1}{2}||x_{A0}-x_{B0}||^{2}$
$\iff||x_{A1}-x_{B1}||^{2}$ $\geq 2||x_{B0}||^{2}+\frac{1}{2}||x_{A0}-x_{B0}||^{2}$
$\iff||x_{A1}||^{2}-||x_{B1}||^{2}$ $\geq\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{A1}^{T}x_{B1}$
This concludes the proof by induction of 7 and 8. Finally, note that if $x_{0}\neq\mathcal{E}$, then either $||x_{A1}-x_{B1}||^{2}>||x_{A0}-x_{B0}||^{2}$, which implies $||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{A1}^{T}x_{B1}$, or $||x_{A2}-x_{B2}||^{2}>||x_{A1}-x_{B1}||^{2}$, which implies $||x_{B2}||^{2}-||x_{A2}||^{2}>\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{A2}^{T}x_{B2}$. Following the same steps of the above proof by induction, it is clear that the strict inequality relation carries through for higher values of $n$, that is for every $n>1$. This concludes the proof of lemma A.2.
We now prove lemma 4.2. Let $x_{0}\in\mathcal{L}$. Fix $n>0$, odd. By the argument of the proof of lemma A.2, it has
$||x_{An}-x_{Bn}||^{2}$ $\geq||x_{Bn-1}||^{2}-||x_{An-1}||^{2}+2x^{T}_{Bn-1}(x_{Bn-1}-x_{An-1})$
$\geq||x_{Bn-1}-x_{An-1}||^{2}+2(||x_{Bn-1}||^{2}-||x_{An-1}||^{2}).$
By lemma A.2, it has
$||x_{An}-x_{Bn}||^{2}$ $\geq||x_{An-1}-x_{Bn-1}||^{2}+2(||x_{Bn-1}||^{2}-||x_{An-1}||^{2})$
$\geq||x_{An-1}-x_{Bn-1}||^{2}+||x_{A0}-x_{B0}||^{2}+4x_{Bn-1}^{T}x_{An-1}$
$\iff||x_{An}-x_{Bn}||^{2}$ $\geq||x_{An-1}+x_{Bn-1}||^{2}+||x_{A0}-x_{B0}||^{2}\geq||x_{A0}-x_{B0}||^{2},$
where lemma A.2 was used in the second inequality. Note that if $x_{0}\neq\mathcal{E}$, our previous results imply that the last three inequalities hold strictly. Next, fix $n>0$, even. By the argument of the proof of lemma A.2, it has
$||x_{An}-x_{Bn}||^{2}$ $\geq||x_{An-1}||^{2}-||x_{Bn-1}||^{2}+2x^{T}_{An-1}(x_{An-1}-x_{Bn-1})$
$\geq||x_{An-1}-x_{Bn-1}||^{2}+2(||x_{An-1}||^{2}-||x_{Bn-1}||^{2}).$
By lemma A.2, it has
$||x_{An}-x_{Bn}||^{2}$ $\geq||x_{An-1}-x_{Bn-1}||^{2}+2(||x_{An-1}||^{2}-||x_{Bn-1}||^{2})$
$\geq||x_{An-1}-x_{Bn-1}||^{2}+||x_{A0}-x_{B0}||^{2}+4x_{An-1}^{T}x_{Bn-1}$
$\iff||x_{An}-x_{Bn}||^{2}$ $\geq||x_{An-1}+x_{Bn-1}||^{2}+||x_{A0}-x_{B0}||^{2}\geq||x_{A0}-x_{B0}||^{2},$
where lemma A.2 was used in the second inequality. Note that if $x_{0}\neq\mathcal{E}$, our previous results imply that the last three inequalities hold strictly. This concludes the proof of lemma 4.2. $\blacksquare$
Proof of Proposition 4.1
First, we prove that the set of Nash equilibria $\mathcal{E}$ is non-empty and finite. Existence of the Nash equilibrium follows from the following two observations: (i) the game admits a potential function; and (ii) the set of best responses to pure strategies is essentially finite.
To see that (i) is true, note that an exact potential for the game is $P:\mathbb{R}^{2K}\to\mathbb{R}$ with
$P(x_{A},x_{B})=V_{A}(x_{A},x_{B})-\frac{||x_{B}||^{2}}{\phi}.$
Second, note that the set of best responses to pure strategies is a subset of
$\mathcal{L}^{p}_{\mathcal{R}}\coloneq\left\{x\in\mathbb{R}^{K}:\exists R\in\mathcal{R}(I),x\in\{\bar{x}(R),\underaccent{\bar}{x}(R)\}\right\},$
which is non-empty and finite because $|I|<\infty$. Hence, for each player, only a finite number of actions are ever best responses to pure strategies. The observation implies that the set of actions that are played in any given best response sequence is finite. We then apply the following well-known results from the literature on potential games. First, the existence of a potential function implies that if the best response sequence converges, it converges to a pure strategy Nash equilibrium. Second, since the number of actions that are ever best responses is finite – and therefore the set of action profiles that are played in a given best response sequence is finite – any best response sequence must converge: the potential function cannot decrease along the best response path, hence the sequence cannot cycle back to a strategy profile that was played and then abandoned in earlier stages of the sequence. This proves that $\mathcal{E}\neq\emptyset$.
Next, we prove that $\mathcal{E}^{\star}$ is non-empty and finite. It is sufficient to note the non-emptyness and finiteness of $\mathcal{E}$, which follows from the previous step of the proof and from the finiteness of $\mathcal{B}$. Finally, we prove that
$\mathcal{E}^{\star}=\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}.$
Let $\hat{x}\in\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$ and assume per contra that $\hat{x}\notin\mathcal{E}^{\star}$. There are two cases, $\hat{x}\in\mathcal{E}$ and $\hat{x}\notin\mathcal{E}$. If $\hat{x}\in\mathcal{E}$, it means that $\exists x\in\mathcal{E}$ such that $V_{A}(x)>V_{A}(\hat{x})$. Since $\mathcal{E}\subseteq\mathcal{L}$, lemma 4.1 implies $||x_{A}-x_{B}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$, but then $\exists x\in\mathcal{L}$ such that $||x_{A}-x_{B}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$, which is a contradiction. Hence it must be $\hat{x}\notin\mathcal{E}$. But we know that the best response sequence initiated at $\hat{x}$ must converge to some pure strategy equilibrium $x\in\mathcal{E}$. Since $\hat{x}\in\mathcal{L}$ by construction, lemma A.2 implies that, according to the best response dynamics initiated at $\hat{x}$, $||x_{An}-x_{Bn}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$ for each $n>1$. Hence it must be that $||x_{A}-x_{B}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$. Since $\mathcal{E}\subseteq\mathcal{L}$, $x\in\mathcal{L}$. But this means that $\exists x\in\mathcal{L}$ such that $||x_{A}-x_{B}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$, which is a contradiction. Hence $\hat{x}\in\mathcal{E}^{\star}$, and $\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}\subseteq\mathcal{E}^{\star}$.
Let $\hat{x}\in\mathcal{E}^{\star}$ and assume per contra that $\hat{x}\notin\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$. Then there exists $x=(x_{A},x_{B})\in\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$ such that $||x_{A}-x_{B}||^{2}>||\hat{x}_{A}-\hat{x}_{B}||^{2}$. Since $\hat{x}\in\mathcal{E}^{\star}$, $x\notin\mathcal{E}$. But the best response sequence initiated at $x$ converges to some pure strategy equilibrium $x^{\prime}\in\mathcal{E}$. Since $x\in\mathcal{L}$ by construction, lemma A.2 implies that, if the best response sequence is initiated at $x$, $||x_{An}-x_{Bn}||^{2}>||x_{A}-x_{B}||^{2}$ for each $n>1$. Hence it must be that $||x^{\prime}_{A}-x^{\prime}_{B}||^{2}>||x_{A}-x_{B}||^{2}$. Since $\mathcal{E}\subseteq\mathcal{L}$, $\exists x^{\prime}\in\mathcal{L}$ such that $||x^{\prime}_{A}-x^{\prime}_{B}||^{2}>||x_{A}-x_{B}||^{2}$, which is a contradiction. Hence $\hat{x}\in\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$, and $\mathcal{E}^{\star}\subseteq\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$. This proves $\mathcal{E}^{\star}=\arg\max_{(x_{A},x_{B})\in\mathcal{L}}||x_{A}-x_{B}||^{2}$. $\blacksquare$
Proof of Proposition 4.2
To prove both point (i) and point (ii) of the proposition it is sufficient to take the derivative of $\hat{V}^{R}_{p}$ with respect of $x^{k}_{R_{i}}$. Fix $(\mathbf{x},\mathbf{s})$ and $k\leq K$. Let $x\in\mathcal{L}(\mathbf{x},\mathbf{s})$ and let $R\in\mathcal{r}(x)$. Note that, for $i\in I$
$\frac{\partial\hat{V}_{p}^{R}(\mathbf{x},\mathbf{s})}{\partial x^{k}_{R_{i}}}=\frac{\bar{w}_{R_{i}}-\underaccent{\bar}{w}_{R_{i}}}{\phi}\left(\bar{x}^{k}(R)-\underaccent{\bar}{x}^{k}(R)\right).$
The derivative is well defined because the ranking induced by $x$ on $(\Delta_{i}(x))_{i\in I}$ is strict as proven in lemma A.1. Next, note that if $i>m_{R}$ we have $\bar{w}_{R_{i}}>\underaccent{\bar}{w}_{R_{i}}$, and therefore
$\text{sign}\frac{\partial\hat{V}^{p}_{R}(\mathbf{x},\mathbf{s})}{\partial x^{k}_{R_{i}}}=\text{sign}\left(\bar{x}^{k}(R)-\underaccent{\bar}{x}^{k}(R)\right),$
while if $i<m_{R}$ we have $\bar{w}_{R_{i}}<\underaccent{\bar}{w}_{R_{i}}$, and therefore
$\text{sign}\frac{\partial\hat{V}^{p}_{R}(\mathbf{x},\mathbf{s})}{\partial x^{k}_{R_{i}}}=\text{sign}\left(\underaccent{\bar}{x}^{k}(R)-\bar{x}^{k}(R)\right).$
which proves part (i). As for part (ii), note that, from the above, it has that,
$\left[\frac{\partial\hat{V}_{p}^{R}(\mathbf{x},\mathbf{s})}{\partial x^{1}_{R_{i}}},\frac{\partial\hat{V}_{p}^{R}(\mathbf{x},\mathbf{s})}{\partial x^{2}_{R_{i}}},...,\frac{\partial\hat{V}_{p}^{R}(\mathbf{x},\mathbf{s})}{\partial x^{K}_{R_{i}}}\right]=$
$\frac{\bar{w}_{R_{i}}-\underaccent{\bar}{w}_{R_{i}}}{\phi}\left[\bar{x}^{1}(R)-\underaccent{\bar}{x}^{1}(R),\bar{x}^{2}(R)-\underaccent{\bar}{x}^{2}(R),...,\bar{x}^{K}(R)-\underaccent{\bar}{x}^{K}(R)\right]$
so that the direction of fastest payoff improvement is $\bar{x}(R)-\underaccent{\bar}{x}(R)$ if $i>m_{R}$ and $\underaccent{\bar}{x}(R)-\bar{x}(R)$ if $i<m_{R}$.
Proof of Proposition 4.3
Consider two distributions, $(\mathbf{x},\mathbf{s}),(\mathbf{\tilde{x}},\mathbf{\tilde{s}})\in\mathcal{V}$ and let $\hat{x}\in\mathcal{E}^{\star}(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$. Note that the payoff $\hat{V}_{p}(\mathbf{x},\mathbf{s})$ does not depend on the medians (means) of the associated symmetric marginal distributions $(\mathbf{x}^{k},\mathbf{s})_{k\leq K}$, so, without loss of generality, we can assume that both $(\mathbf{x},\mathbf{s})$ and $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$ have marginal distributions that are centered around 00. We want to show that if $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$ is a spread of $(\mathbf{x},\mathbf{s})$ in direction $\hat{x}_{A}-\hat{x}_{B}$, then $\hat{V}_{p}(\tilde{\mathbf{x}},\tilde{\mathbf{s}})>\hat{V}_{p}(\mathbf{x},\mathbf{s})$.
Assume that $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$ is a spread of $(\mathbf{x},\mathbf{s})$. Let $\tilde{V}_{p}(\hat{x})$ be parties’ expected payoff when they play $\hat{x}$ and the voter distribution is $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$, and let $V_{p}(\hat{x})$ be parties’ expected payoff when they play $\hat{x}$ and the voter distribution is $(\mathbf{x},\mathbf{s})$. We proceed in two steps. First, we show that $\tilde{V}_{p}(\hat{x})>V_{p}(\hat{x})=\frac{\nu(1)+\nu(0)}{2}+\frac{2}{\phi}||\hat{x}_{A}||^{2}$. Starting from this fact, we can consider the best-response dynamics based on distribution $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$ and use the same reasoning as in the proof of lemma 4.2 to show that $\hat{V}_{p}(\tilde{\mathbf{x}},\tilde{\mathbf{s}})\geq\tilde{V}_{p}(\hat{x})$. For the first step, note that
$\tilde{V}_{p}(\hat{x})=\frac{\nu(1)+\nu(0)}{2}+\frac{2}{\phi}\sum_{i\in\tilde{I}}\bar{w}_{\tilde{R}_{i}}\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A},$
where $\tilde{I}$ indexes types in $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$, $\tilde{R}\in\mathcal{R}(\tilde{I})$, is such that $\tilde{R}\in\mathcal{r}(\hat{x})$, and the weights $(\bar{w}_{\tilde{R}_{i}})_{i\in\tilde{I}}$ are constructed as usual, based on order $\tilde{R}$ and distribution $(\mathbf{\tilde{x}},\mathbf{\tilde{s}})$.
Similarly, it has
$V_{p}(\hat{x})=\frac{\nu(1)+\nu(0)}{2}+\frac{2}{\phi}\sum_{i\in I}\bar{w}_{R_{i}}x_{R_{i}}^{T}\hat{x}_{A}$
where $I$ indexes types in $(\mathbf{x},\mathbf{s})$, $R\in\mathcal{R}(I)$, is such that $R\in\mathcal{r}(\hat{x})$, and the weights $(\bar{w}_{R_{i}})_{i\in I}$ are based on order $R$ and distribution $(\mathbf{x},\mathbf{s})$.
In order to prove that $\sum_{i\in\tilde{I}}\bar{w}_{\tilde{R}_{i}}\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}>\sum_{i\in I}\bar{w}_{R_{i}}x_{R_{i}}^{T}\hat{x}_{A}$ it is sufficient to follow the same steps of used in proposition 3.2 to show that $\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}>\bar{x}-\underaccent{\bar}{x}$. In particular, the proof follows from the one of proposition 3.2 if we impose (i) $I^{\prime}=\tilde{I}$; (ii) for each $i\in\tilde{I}$, $\bar{w}^{\prime}_{i}=\underaccent{\bar}{w}_{N-i}^{\prime}=\bar{w}_{\tilde{R}_{i}}$, $x^{\prime}_{i}=-x^{\prime}_{N-i}=\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}$, $s^{\prime}_{i}=s^{\prime}_{N-i}=\tilde{s}_{\tilde{R}_{i}}$; and (iii) for each $i\in I$, $\bar{w}_{i}=\underaccent{\bar}{w}_{N-i}=\bar{w}_{R_{i}}$, $x_{i}=-x_{N-i}=x_{R_{i}}^{T}\hat{x}_{A}$, and $s_{i}=s_{N-i}=s_{R_{i}}$.
As for the second part of the proof, we start by noting that the inductive step of the proof of lemma A.2 still holds at when we consider distribution $(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$ and initiate the best response dynamics at $\hat{x}$: by using the exact same steps, one can indeed show that if the inequality 7 holds at some odd step $n$ of the best response sequence, then it must hold at step $n+2$, and 8 must hold at $n+1$ of the best response sequence. Hence, it is sufficient to show that $||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||\hat{x}_{A}-\hat{x}_{B}||^{2}+2x_{A1}^{T}x_{B1}$, that is, that inequality 7 indeed holds at $n=1$.
Note that
$||x_{A1}-x_{B1}||^{2}\geq 2\phi\left[\tilde{V}_{p}(\hat{x}_{A},\hat{x}_{B})-\frac{\nu(1)+\nu(0)}{2}\right]>$
$>2\phi\left[V_{p}(\hat{x}_{A},\hat{x}_{B})-\frac{\nu(1)+\nu(0)}{2}\right]>||\hat{x}_{A}-\hat{x}_{B}||^{2}$
So that
$||x_{A1}-x_{B1}||^{2}>||x_{A1}-x_{B1}||^{2}=||x_{A0}-x_{B0}||^{2}$
which, following the same steps of the proof of lemma A.2, implies
$||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||x_{A0}-x_{B0}||^{2}+2x_{A1}^{T}x_{B1}.$ (9)
This proves that the result of lemma A.2 holds when the best response dynamics is initiated at $\hat{x}$. Next note that, as shown in the proof od lemma A.2, inequality 9 implies that
$||x_{An}-x_{Bn}||^{2}>||x_{A0}-x_{B0}||^{2}=||\hat{x}_{A}-\hat{x}_{B}||^{2},\quad n=1,2,3,...$
so that the best response dynamics initiated after the spread will keep parties further apart from each other than before the spread. But by proposition 4.1, the best response dynamics will converge to some Nash equilibrium, $x^{\star}\in\mathcal{E}(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$. And since parties will be strictly further apart in such a Nash equilibrium than under $\hat{x}$, it must be that that $\tilde{V}_{p}(x^{\star})>\hat{V}_{p}(\mathbf{x},\mathbf{s})$. But this implies $\hat{V}_{p}(\tilde{\mathbf{x}},\tilde{\mathbf{s}})>\hat{V}_{p}(\mathbf{x},\mathbf{s})$, which proves the statement. $\blacksquare$

### v2

@@
 Our main theoretical results highlight a novel incentive to exacerbate societal divisions.
-In the conclusive sections of the paper we discuss examples of how this might occur, relating the main themes of the previous sections to the growing literature on zero-sum thinking (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), common interest vs conflict of interest policy issues (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023), and identity politics (e.g.
+In the conclusive sections of the paper we discuss examples of how this might occur, relating the main themes of the previous sections to the growing literature on zero-sum thinking (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023; Ali et al.,, 2024), common interest vs conflict of interest policy issues (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023), and identity politics (e.g.
 Shayo,, 2009; Bonomi et al.,, 2021; Gennaioli and Tabellini,, 2023).
@@
 [^4] Indeed, if voters had heterogeneous priors or idiosyncratic private information, public information about common interest issues would bring them closer together, decreasing the equilibrium expected payoff of politicians.
-Our last application concerns party identity, modeled building on (Bonomi et al.,, 2021).
+Our last application concerns party identity, modeled building on Bonomi et al., (2021).
 We show that prompting voters to identify with the party they are more likely to vote for is not only valuable but also “efficient” for politicians, in that it distorts all voters’ bliss points in the direction of the fastest payoff improvement for candidates.
@@
 The shock has a continuous cumulative distribution $F$, and captures random events that influence voters’ relative assessment of party $B$, including, for instance, new information about the state of the economy, political scandals, or the effect of a candidate’s personality traits and characteristics on their appeal to voters.
-We assume that voters vote sincerely, taking into account $x_{A}$, $x_{B}$, and the popularity shock $\epsilon$.
+We assume that voters vote sincerely, taking into account $x_{A}$, $x_{B}$, and the popularity shock $\epsilon$ [^7] Note that sincere voting is without loss of generality with atomistic voters, a scenario allowed in our setup.
+In fact, we only assume a finite number of voter types.
 In particular, for each $x_{A},x_{B}$, let $\Delta_{i}(x_{A},x_{B})=v(x_{A},x_{i})-v(x_{B},x_{i})$.
@@
 Part (i) requires that rents increase in the number of votes.
-[^8] Proposition 2.1 would generalize to the case where $R$ is weakly increasing on $[0,\frac{1}{2}]$ and not constant on $[0,\frac{1}{2})$.
+[^9] Proposition 2.1 would generalize if $R$ was weakly increasing on $[0,\frac{1}{2}]$ and not constant on $[0,\frac{1}{2})$.
 In that case, party divergence would obtain for some distributions of voter types.
@@
 An alternative interpretation is that risk-neutral politicians might use the rents to maintain their support through transfers that increase their constituents’ (concave) consumption utility, or by investments in a propaganda technology with convex costs.
-Let $\nu:[0,1]\to\mathbb{R}$, $\nu=U\circ R$, be the function mapping a party’s vote share to its rents, and without any loss of generality, let payoffs be normalized so that $\nu(1)-\nu(0)=1$.
+Let $\nu:[0,1]\to\mathbb{R}$, $\nu=U\circ R$, be the function mapping a party’s vote share to its payoff, and without any loss of generality, let payoffs be normalized so that $\nu(1)-\nu(0)=1$.
 The following holds when assumptions 1 and 2 are combined.
@@
 Proposition 2.1 (Party Divergence)
-In every pure strategy equilibrium, candidates offer different policy platforms.
+If a pure strategy equilibrium exists, candidates offer different policy platforms.
 The intuition for the result is as follows.
@@
 [^13] In this case, dimension indexes $k$ are redundant, and omitted.
-Let $\epsilon\sim U(-\phi,\phi$), so that the shock is uniformly distributed, and assume that $h^{k}(|x-x_{i}|)=-\alpha(x^{k}-x^{k}_{i})^{2}$, where $\alpha>0$ captures how important policies are for individual payoffs and, without loss of generality, we set $\alpha=1$.
+Let $\epsilon\sim U(-\phi,\phi$), so that the shock is uniformly distributed, and assume that $h(|x-x_{i}|)=-\alpha(x-x_{i})^{2}$, where $\alpha>0$ captures how important policies are for individual payoffs and, without loss of generality, we set $\alpha=1$.
 With one policy dimension, we can think of voter heterogeneous preferences as capturing the classical economy-based conflict of interest between right-wing and left-wing voters, as in the following simple example.
@@
 Hence, this example satisfies the assumptions of our model for $x=\tau$, $x_{i}=\tau_{i}$, and $g(x_{i})=\psi-cx_{i}+\frac{c}{2}x_{i}^{2}$.
-First, we show that the game admits a pure strategy equilibrium, and there are exactly two platforms that can be offered in such an equilibrium.
-One policy lies to the left of the risk-neutral equilibrium policy $x^{RN}$.
-The other lies to the right of such policy.
 For the next propositions, assume that voters are ordered so that $x_{1}\leq x_{2}\leq...\leq x_{N}$, and let $x_{m}$ be the median voter type.
 If the median voter does not exist, define $x_{m}$ to be the average of the two “most median” voter types.
-Proposition 3.1 (Unidimensional Equilibrium)
-The game admits a pure strategy equilibrium.
-Moreover, there exist two platforms, $\underaccent{\bar}{x},\bar{x}\in\mathbb{R}$ with $x_{1}<\underaccent{\bar}{x}<\bar{x}<x_{N}$, such that, in every pure strategy equilibrium, one candidate offers $\underaccent{\bar}{x}$ while the other candidate offers $\bar{x}$.
+We show that the game admits a pure strategy equilibrium, and there are exactly two platforms that can be offered in such an equilibrium.
+One policy lies to the left of the risk-neutral equilibrium policy $x^{RN}$, where
+$x^{RN}=\sum_{i=1}^{N}\frac{R(s_{i}+\sum_{j<i}s_{j})-R(\sum_{j<i}s_{j})}{\bar{R}}x_{i},$
+with $s_{0}=0$. The other lies to the right of such policy.
+Proposition 3.1 (Equilibrium)
+The game has a pure strategy equilibrium. In particular:
+  1. (i)
+    There exist two platforms, $\underaccent{\bar}{x},\bar{x}\in\mathbb{R}$ with $x_{1}<\underaccent{\bar}{x}<x^{RN}<\bar{x}<x_{N}$, such that, in every pure strategy equilibrium, one candidate offers $\underaccent{\bar}{x}$ and the other offers $\bar{x}$.
 It has
     $\underaccent{\bar}{x}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i},\quad\quad\quad\quad\bar{x}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$
-where $w_{i},\omega_{i}\in(0,1)$, $\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=\sum_{i=1}^{N}\bar{w}_{i}=1$, and $\bar{w}_{i}>(<)\text{ }\underaccent{\bar}{w}^{i}$ if $x_{i}>(<)\text{ }x_{m}$, where $x_{m}$ is the median voter type.
-For each $i$, the equilibrium weights $\bar{w}_{i}$ and $\underaccent{\bar}{w}_{i}$ that parties attribute to voter group $i$ are proportional to their payoff increase from winning the votes of such group.
+    where $\bar{w}_{i},\underaccent{\bar}{w}_{i}\in(0,1)$ for each $i\in I$, $\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=\sum_{i=1}^{N}\bar{w}_{i}=1$.
+  2. (ii)
+    For each $i\in I$, the equilibrium weights $\bar{w}_{i}$ and $\underaccent{\bar}{w}_{i}$ attributed to $x_{i}$ are proportional to their payoff increase from winning the votes of group $i$.
+In particular
+    $\underaccent{\bar}{w}_{i}=\frac{\nu(\underaccent{\bar}{s}_{i})-\nu(\underaccent{\bar}{s}_{i-1})}{\nu(1)-\nu(0)},\quad\quad\quad\quad\bar{w}_{i}=\frac{\nu(\bar{s}_{i+1})-\nu(\bar{s}_{i})}{\nu(1)-\nu(0)},$
+    where, for each $i\in I$, $\underaccent{\bar}{s}_{i}=\sum_{j\leq i}s_{j}$ and $\bar{s}_{i}=\sum_{j\geq i}s_{j}$.
+  3. (iii)
+    One candidate attributes more weight to the policy demands of those to the left of the median voter, the other candidate attributes more weight to the demands of those to the right of the median voter.
+In particular, for each $i\in I$,
+    $x_{i}>(<)\text{ }x_{m}\implies\bar{w}_{i}>(<)\text{ }\underaccent{\bar}{w}^{i}$
+    where $x_{m}$ is the median voter type.
+  4. (iv)
     Finally, the equilibrium expected payoff of each party is proportional to the (squared) distance between the two equilibrium platforms $(\bar{x}-\underaccent{\bar}{x})^{2}$, up to a constant term.
@@
 We say that candidate $p\in\{A,B\}$ benefits from attracting voter group $i\in I$ if $\frac{\partial\pi^{p}_{i}}{\partial x_{i}}\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}>0$, while we say that she benefits from alienating voter group $i\in I$ if $\frac{\partial\pi^{p}_{i}}{\partial x_{i}}\frac{\partial\hat{V}_{p}(\mathbf{x},\mathbf{s})}{\partial x_{i}}<0$.
-In other words, a right-wing candidate benefits from attracting voter type $i$ if a change in $i$’s views on redistribution pushes her to support the political right and also increases the equilibrium payoff of the right-wing party.
-Conversely, the right-wing party benefits from alienating voter type $i$ if an increase in $i$’s support for redistribution makes the right-wing party better off in equilibrium, despite increasing left-wing support from group $i$.
-At this point, one could be skeptical that a candidate ever benefits from alienating a group of voters, even if it means that the rest of the voter distribution remains unchanged and that group of voters will be less likely to support the party in the new equilibrium.
+In other words, we say that candidate $A$ benefits from attracting voter group $i$ if a change in voter group $i$’s views that increase its support for candidate $A$ improves candidate $A$’s equilibrium payoff.
+Conversely, candidate A benefits from alienating voting group $i$ if an increase in group $i$’s support for party $A$ decreases party $A$’s equilibrium payoff.
+At this point, one could be skeptical that a candidate ever benefits from alienating a group of voters, given that it means that the rest of the voter distribution remains unchanged and that group of voters will be less likely to support the party in the new equilibrium.
 Surprisingly, our next result suggests that each party always has the incentive to alienate at least one voter group: the most extremist supporters of the rival candidate.
@@
 Strict concavity guarantees that party best responses never fall at points where payoffs are non-differentiable, which greatly simplifies the analysis.
-[^17] The assumption can be relaxed and, in particular, we can allow for a jump in payoff when parties reach the majority of votes.
+[^18] The assumption can be relaxed.
+In particular, we can allow for a jump in payoff when parties reach the majority of votes, provided that a median voter type exists.
 Second, for some of the main results of the section we restrict the attention to symmetric voter distribution, which we prove to admit a pure strategy Nash equilibrium.
-[^18] In the more general formulation of the game, the existence of a mixed strategy equilibrium is guaranteed by the Glicksberg theorem, but the existence of a pure strategy equilibrium is not, because the utility function of a party is not quasi-concave in that party’s strategy.
+[^19] In the more general formulation of the game, the existence of a mixed strategy equilibrium is guaranteed by Glicksberg’s existence result, but the existence of a pure strategy equilibrium is not, because the utility function of a party is not quasi-concave in that party’s strategy.
 Definition 3
@@
 The equilibrium analysis presents some technical challenges.
-First, parties’ expected payoffs are not differentiable at platform profiles $(x_{A},x_{B})$ such that some voter group $i$ is indifferent between voting for party $A$ or party $B$.
+First, parties’ expected payoffs are not differentiable at platform profiles $(x_{A},x_{B})$ such that two voter groups, $i$ and $j$, have the same relative support for party $A$ (i.e., $\Delta_{i}(x_{A},x_{B})=\Delta_{j}(x_{A},x_{B})$).
 Second, a party’s expected payoff is not quasi-concave in that party platform – best responses are not continuous in the competitors’ platforms.
@@
 Panel (b) plots a symmetric bidimensional voter distribution and the associated local equilibria.
-In both panels: (i) ”x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) two circles of the same color represent a pair of local equilibrium platforms; (iii) we assume that party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$.
+In both panels: (i) “x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) two circles of the same color represent a pair of local equilibrium platforms; (iii) we assume that party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$.
 Figure 1 panel (a) reports a voter distribution over a bidimensional policy space, as well as the set of local equilibria, found applying the algorithm outlined in the previous paragraph.
-Lemma 4.1 tells us that among the existing local equilibria, the one identified by the two green dots is associated to the highest party expected payoff since the platforms proposed by the two politicians differ the most.
+Lemma 4.1 tells us that among the existing local equilibria, the one identified by the two dark dots is associated to the highest party expected payoff since the platforms proposed by the two politicians differ the most.
 In panel (b) we report the voter types and local equilibria for a symmetric voter distribution.
-In this case, the local equilibrium in blue is also the Nash equilibrium where parties expect to benefit the most, as per the results of proposition 4.1.
+In this case, the local equilibrium identified by the two dark dots is also the Nash equilibrium where parties expect to benefit the most, as per the results of proposition 4.1.
 In section 3.1 we found that, with one policy dimension, politicians benefit from pushing voters to the left of the median voter further to the left, and those to the right further to the right.
@@
 Figure 2 reports the direction in which voter opinions should change for parties’ equilibrium payoffs to increase the fastest, for the two local equilibria of figure 1, panel (a).
-The red arrows represent this direction for the best local equilibrium (identified by the green circles), while the blue arrows identify the direction for the other local equilibrium (identified by the white circles).
+The red arrows represent this direction for the best local equilibrium (identified by the dark circles), while the blue arrows identify the direction for the other local equilibrium (identified by the white circles).
 In both cases, the change makes supporters of party $A$ on average more different from supporters of party $B$.
@@
 In the next section, we interpret themes on party identity, zero-sum thinking, and common interest vs conflict of interest policies in light of these incentives.
-5 Applications
+5 Extensions and Applications
 In this section, we briefly discuss applications of our theory to zero-sum policies, parties’ information provision incentives, and party identity.
@@
 A recent stream of literature in political science and economics has stressed how politicians tend to focus on divisive, zero-sum, issues over common interest issues, even if the latter entail policies that could benefit a larger share of voters (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023).
-This phenomenon is potentially amplified by the fact that, on the one hand, voters seem to be prone to social and political “zero-sum thinking,” the perception that gains from one social group must necessarily come at the expense of another (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), and often lack the information to assess the “common value” nature of policies (Matsusaka and Kendall,, 2022), while politicians’ rhetoric often paints politics in a zero-sum fashion.
+This phenomenon is potentially amplified by the fact that, on the one hand, voters seem to be prone to social and political “zero-sum thinking,” the perception that gains from one social group must necessarily come at the expense of another (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023; Ali et al.,, 2024), and often lack the information to assess the “common value” nature of policies (Matsusaka and Kendall,, 2022), while politicians’ rhetoric often paints politics in a zero-sum fashion.
 Our simple model can provide a new perspective on why politicians might have an incentive to focus on more divisive issues, stress the zero-sum component of policy-making, and keep voters informed on zero-sum issues and uninformed about common interest policies.
@@
 $\lim_{R_{m}\rightarrow\bar{R}^{-}}\bar{w}_{m}=\lim_{R_{m}\rightarrow\bar{R}^{-}}\underaccent{\bar}{w}_{m}=1$
-so that, when $R\to\bar{R}$, $\bar{x}\to x^{o}$ and $\underaccent{\bar}{x}\to x^{o}$. This result is captured by the following remark.
+so that, when $R_{m}\to\bar{R}$, $\bar{x}\to x^{o}$ and $\underaccent{\bar}{x}\to x^{o}$. This result is captured by the following remark.
 Remark 11
@@
 Proof of Proposition 3.1
-By proposition 2.1, we know that for any platform $x\in\mathbb{R}$, $x$ is never the best response to $x$.
+We start from part (i).
 Fix a distribution $(\mathbf{x},\mathbf{s})$ and order $i,...,N$ so that $x_{1}\leq x_{2}\leq...\leq x_{N}$.
@@
 As a first step we show that, if $x_{p}$ is a best response to $x_{-p}$, then either $x_{p}>x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$, for $\bar{w}_{i}=\frac{\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})}{\nu(1)-\nu(0)}$, or $x_{p}<x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$, for $\underaccent{\bar}{w}_{i}=\frac{\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})}{\nu(1)-\nu(0)}$.
+By proposition 2.1, we know that for any platform $x\in\mathbb{R}$, $x$ is never the best response to $x$.
 Assume that $x_{p}$ is a best response to $x_{-p}$.
@@
 Hence, the game admits at most two best responses, $\underaccent{\bar}{x}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$ and $\bar{x}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$.
-In order to show that a pure strategy Nash equilibrium exists, we need to show that $\bar{x}$ and $\underaccent{\bar}{x}$ are the best responses to each other.
-It suffices to show that $\underaccent{\bar}{x}<\bar{x}$, as we argue below.
-First, by proposition 2.1, $\underaccent{\bar}{x}$ is not the best response to $\underaccent{\bar}{x}$ and $\bar{x}$ is not the best response to $\bar{x}$.
-Hence, by the previous step of the proof, $\underaccent{\bar}{x}$ is the only possible best response to $\bar{x}$ and $\bar{x}$ is the only possible best response to $\underaccent{\bar}{x}$.
-Second, note that, for each $p=A,B$, $V_{p}$ is continuous, increasing in $x_{p}$ when $x_{p}<\underaccent{\bar}{x}$, decreasing in $x_{p}$ when $x_{p}>\bar{x}$.
-Hence, given the strict concavity of $V_{p}$ in $x_{p}$ when $x_{p}\neq x_{-p}$, $x_{p}$ is the best response to $\underaccent{\bar}{x}$ if $x_{p}>\underaccent{\bar}{x}$ and it satisfies the FOC of the respective maximization problem.
-The FOC requires $x_{p}=\bar{x}$.
-Similarly, $x_{p}$ is the best response to $\bar{x}$ if $x_{p}<\bar{x}$ and it satisfies the FOC of the respective maximization problem.
-The FOC requires $x_{p}=\underaccent{\bar}{x}$.
-Hence, $\underaccent{\bar}{x}<\bar{x}$ is a sufficient condition for $\bar{x}$ and $\underaccent{\bar}{x}$ to be the best responses to each other.
-We now show $\underaccent{\bar}{x}<\bar{x}$.
+Next, we show that for each $p\in\{A,B\}$ and each $x_{B}\in\mathbb{R}$, $A$ has at least one best response to any $x_{B}$.
+Fix $x_{B}\in\mathbb{R}$.
+First, note that any platform choice outside $[x_{1},x_{N}]$ is strictly dominated by either $x_{1}$ or $x_{N}$.
+We can then focus on looking for best responses in $[x_{1},x_{N}]$.
+Next, note that $V_{A}(\cdot,x_{B})$ is continuous on $[x_{1},x_{N}]$.
+Hence, by the extreme value theorem, for any $x_{B}$, $A$ has at least one best response to $x_{B}$.
+For the same reason, $B$ has at least one best response to any $x_{A}\in\mathbb{R}$.
+Given that the best response correspondence is never empty valued, to show that a pure strategy Nash equilibrium exists it suffices to show that $\underaccent{\bar}{x}<\bar{x}$.
 First, note that since $\underaccent{\bar}{w}_{i},\bar{w}_{i}\in[0,1]$ for $i=1,...,N$, and $\sum_{i=1}^{N}\bar{w}_{i}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=1$, $(\bar{w}_{i},x_{i})_{i\in I}$ and $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$ are probability distributions.
@@
 where the inequality hold strictly for every $i>1$.
-Having proved strict first order stochastic dominance, note that $\bar{x}$ is the expected value of $(\bar{w}_{i},x_{i})_{i\in I}$ and $\underaccent{\bar}{x}$ is the expected value of $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$.
+Finally, note that $\bar{x}$ is the expected value of $(\bar{w}_{i},x_{i})_{i\in I}$ and $\underaccent{\bar}{x}$ is the expected value of $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$.
 Strict first order stochastic dominance implies $\bar{x}>\underaccent{\bar}{x}$.
 This concludes the proof that the pure strategy equilibrium exists, and that only $\bar{x}$ and $\underaccent{\bar}{x}$ will be played in such equilibrium.
-Next, note that for each $i\in I$, it has $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$.
+To prove part (ii), note that for each $i\in I$, it has $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$.
 The former weight represents the payoff gain of the party offering $\bar{x}$ when she wins type $i$.
@@
 Using the same argument, it is easy to show that $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$ is the payoff gain of the party offering $\underaccent{\bar}{x}$ from winning voter group $i$.
-Note that, by remark 1, it has
+Part (iii) follows immediately. Note that, by remark 1, it has
 $\bar{s}_{i}<\frac{1}{2}\implies\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})>\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$
@@
 In order to prove that $\sum_{i\in\tilde{I}}\bar{w}_{\tilde{R}_{i}}\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}>\sum_{i\in I}\bar{w}_{R_{i}}x_{R_{i}}^{T}\hat{x}_{A}$ it is sufficient to follow the same steps of used in proposition 3.2 to show that $\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}>\bar{x}-\underaccent{\bar}{x}$.
-In particular, the proof follows from the one of proposition 3.2 if we impose (i) $I^{\prime}=\tilde{I}$; (ii) for each $i\in\tilde{I}$, $\bar{w}^{\prime}_{i}=\underaccent{\bar}{w}_{N-i}^{\prime}=\bar{w}_{\tilde{R}_{i}}$, $x^{\prime}_{i}=-x^{\prime}_{N-i}=\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}$, $s^{\prime}_{i}=s^{\prime}_{N-i}=\tilde{s}_{\tilde{R}_{i}}$; and (iii) for each $i\in I$, $\bar{w}_{i}=\underaccent{\bar}{w}_{N-i}=\bar{w}_{R_{i}}$, $x_{i}=-x_{N-i}=x_{R_{i}}^{T}\hat{x}_{A}$, and $s_{i}=s_{N-i}=s_{R_{i}}$.
-As for the second part of the proof, we start by noting that the inductive step of the proof of lemma A.2 still holds at when we consider distribution $(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$ and initiate the best response dynamics at $\hat{x}$: by using the exact same steps, one can indeed show that if the inequality 7 holds at some odd step $n$ of the best response sequence, then it must hold at step $n+2$, and 8 must hold at $n+1$ of the best response sequence.
+In particular, the proof follows from the one of proposition 3.2 if we impose (i) $I^{\prime}=\tilde{I}$; (ii) for each $i\in\tilde{I}$, $\bar{w}^{\prime}_{i}=\underaccent{\bar}{w}_{N+1-i}^{\prime}=\bar{w}_{\tilde{R}_{i}}$, $x^{\prime}_{i}=-x^{\prime}_{N+1-i}=\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}$, $s^{\prime}_{i}=s^{\prime}_{N+1-i}=\tilde{s}_{\tilde{R}_{i}}$; and (iii) for each $i\in I$, $\bar{w}_{i}=\underaccent{\bar}{w}_{N+1-i}=\bar{w}_{R_{i}}$, $x_{i}=-x_{N+1-i}=x_{R_{i}}^{T}\hat{x}_{A}$, and $s_{i}=s_{N+1-i}=s_{R_{i}}$.
+As for the second part of the proof, we start by noting that the inductive step of the proof of lemma A.2 still holds when we consider distribution $(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$ and initiate the best response dynamics at $\hat{x}$: by using the exact same steps, one can indeed show that if the inequality 7 holds at some odd step $n$ of the best response sequence, then it must hold at step $n+2$, and 8 must hold at $n+1$ of the best response sequence.
 Hence, it is sufficient to show that $||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||\hat{x}_{A}-\hat{x}_{B}||^{2}+2x_{A1}^{T}x_{B1}$, that is, that inequality 7 indeed holds at $n=1$.